### v1

# Bootstrapping Estimators based on the Block Maxima Method

Abstract
The block maxima method is a standard approach for analyzing the extremal behavior of a potentially multivariate time series. It has recently been found that the classical approach based on disjoint block maxima may be universally improved by considering sliding block maxima instead. However, the asymptotic variance formula for estimators based on sliding block maxima involves an integral over the covariance of a certain family of multivariate extreme value distributions, which makes its estimation, and inference in general, an intricate problem. As an alternative, one may rely on bootstrap approximations: we show that naive block-bootstrap approaches from time series analysis are inconsistent even in i.i.d. situations, and provide a consistent alternative based on resampling circular block maxima. As a by-product, we show consistency of the classical resampling bootstrap for disjoint block maxima, and that estimators based on circular block maxima have the same asymptotic variance as their sliding block maxima counterparts. The finite sample properties are illustrated by Monte Carlo experiments, and the methods are demonstrated by a case study of precipitation extremes.
Keywords. Bootstrap Consistency; Disjoint and Sliding Block Maxima; Extreme Value Statistics; Pseudo Maximum Likelihood Estimation; Time Series Analysis.
MSC subject classifications. Primary 62F40, 62G32; Secondary 62E20.
1 Introduction
Let $(\bm{X}_{t})_{t\in\mathbb{Z}}$ with $\bm{X}_{t}=(X_{t1},\dots,X_{td})^{\top}$ denote a stationary time series in $\mathbb{R}^{d}$. For block size $r\in\mathbb{N}$, let $\bm{M}_{r}=\max_{t=1}^{r}\bm{X}_{t}=(\max_{t=1}^{r}X_{t1},\dots,\max_{t=1}^{r}X_{t1})^{\top}$ denote the componentwise maximum over $r$ successive observations. In view of a version of the classical extremal types theorem for multivariate time series, the distribution of $\bm{M}_{r}$ may be approximated by a multivariate extreme value distribution, for sufficiently large block size $r$ (see Section 2.1 for details). Both the distribution of $\bm{M}_{r}$ and parameters related to its potential weak limit distribution are common target parameters in extreme value statistics, with applications in finance, insurance or environmental statistics among others (Beirlant et al., , 2004; Katz et al., , 2002; Philip et al., , 2020).
Suppose $\bm{X}_{1},\dots,\bm{X}_{n}$ is an observed stretch from $(\bm{X}_{t})_{t\in\mathbb{Z}}$. The classical block maxima method for estimating parameters associated with the law of $\bm{M}_{r}$ (or its limiting law for $r\to\infty$) consists of dividing the sampling period $\{1,\dots,n\}$ into $m$ disjoint blocks of length $r$ (for simplicity, we assume that $n=mr$), and of using the sample of (disjoint) block maxima $(\bm{M}_{1,r}^{({\operatorname{db}})},\dots,\bm{M}_{m,r}^{({\operatorname{db}})})$ as a starting point for statistical methods; here $\bm{M}^{({\operatorname{db}})}_{i,r}=\max_{t=(i-1)r+1,\dots,ir}\bm{X}_{t}$ denotes the $i$th disjoint block maximum. This classical approach can often be improved by instead considering the sample of sliding block maxima $(\bm{M}_{1,r}^{({\operatorname{sb}})},\dots,\bm{M}_{n-r+1,r}^{({\operatorname{sb}})})$ as a starting point; here $\bm{M}^{({\operatorname{sb}})}_{i,r}=\max_{t=i,\dots,i+r-1}\bm{X}_{t}$ denotes the block maximum of size $r$ starting at “day” $i$. Note that this sample also provides a stationary, but strongly auto-correlated sample from the law of $\bm{M}_{r}$. It was shown in Bücher and Segers, 2018a ; Zou et al., (2021) among others that estimators that explicitly or implicitly involve empirical means of the sliding block maxima sample typically have the same expectation but a slightly smaller variance than the respective counterparts based on the disjoint block maxima sample.
It seems natural that a lower estimation variance offers the possibility of constructing smaller confidence intervals for a given confidence level. The main goal of this work is to find universal practical solutions for this heuristic based on suitable bootstrap approximations. A challenge consists of the fact that the large sample asymptotics of the sliding block maxima method depend on a blocking of blocks approach involving some intermediate blocking parameter converging to infinity at a well-balanced rate, whence standard bootstrap approaches for time series like the block bootstrap (Lahiri, , 2003) would depend on such an intermediate blocking parameter as well (which must be avoided due to the typically small effective sample size in extreme value analysis). However, as we will show, a novel approach based on what we call the circular (sliding) block maxima sample allows for consistent distributional approximations without relying on such an intermediate parameter sequence. At the same time, the method is computationally attractive because it avoids recalculation of any block maxima for the bootstrap sample.
In general, various variants of the bootstrap have been routinely applied in extreme value statistics (see, for instance, Eastoe and Tawn, , 2008; Huser and Davison, , 2014). However, to the best of our knowledge, respective statistical theory is only available for the peak-over-threshold method (Peng and Qi, , 2008; Drees, , 2015; Davis et al., , 2018; Kulik and Soulier, , 2020; Jentsch and Kulik, , 2021; de Haan and Zhou, , 2024). Concerning the block maxima method, within the regime where the block size is treated as a parameter sequence $r=r_{n}$ converging to infinity, the only reference we are aware of is de Haan and Zhou, (2024), who derive an asymptotic expansion for the tail quantile process of bootstrapped disjoint block maxima. However, their results require an underlying i.i.d. sequence, they only concern the disjoint block maxima method and they do not eventually imply desirable classical consistency statements. While our main focus is on the sliding block maxima method, we establish consistency results for the disjoint block maxima method as a by-product, for serially dependent data.
Finally, we also derive limit results for empirical means and estimators based on the circular (sliding) block maxima sample, and prove that they show the same favorable asymptotic behavior as the sliding block maxima counterparts. As we will discuss in the conclusion, the result may be of independent interest when dealing with non-stationary situations involving, for example, trends; note that the latter is a typical use case for climate extremes.
The remaining parts of this paper are organized as follows: basic model assumptions and known results on the disjoint and sliding block maxima method are collected in Section 2. The sample of circular block maxima is introduced in Section 3 and supplemented by a central limit theorem. Respective bootstrap schemes are proposed in Section 4, where we also provide respective consistency statements. The previous high-level results are applied to a specific estimation problem for univariate heavy tailed time series in Section 5. Finite-sample results on a large-scale simulation study and a small case study are presented in Sections 6 and 7, respectively. Section 8 concludes, and all proofs and technical conditions are collected in Sections 9-11. Some additional theoretical results and simulation results are deferred to a sequence of appendices.
Throughout, weak convergence of sequences of random vectors/distributions is denoted by $\rightsquigarrow$. The code used for the simulation study is available on Github, see Staud, (2024).
2 Mathematical preliminaries
2.1 Basic model assumptions
An extension of the classical extremal types theorem to strictly stationary time series (Leadbetter, , 1983) implies that, under suitable broad conditions, affinely standardized maxima extracted from a stationary time series converge to the generalized extreme value distribution (GEV). This was generalized to the multivariate case in Hsing, (1989), where the marginals are necessarily GEV-distributed. We make this an assumption, and additionally require the scaling sequences to exhibit some common regularity inspired by the max-domain of attraction condition in the i.i.d. case (de Haan and Ferreira, , 2006).
Condition 2.1 (Multivariate max-domain of attraction).
Let $(\bm{X}_{t})_{t\in\mathbb{Z}}$ denote a strictly stationary time series in $\mathbb{R}^{d}$ with continuous margins. There exist sequences $(\bm{a}_{r})_{r}=(a_{r}^{(1)},\dots,a_{r}^{(d)}))_{r}\subset(0,\infty)^{d},(\bm{b}_{r})_{r}=(b_{r}^{(1)},\dots,b_{r}^{(d)})_{r}\subset\mathbb{R}^{d}$ and $\bm{\gamma}=(\gamma^{(1)},\dots,\gamma^{(d)})\in\mathbb{R}^{d}$, such that, for any $s>0$ and $j\in\{1,\dots,d\}$,
$\lim\limits_{r\to\infty}\frac{a_{\lfloor rs\rfloor}^{(j)}}{a_{r}^{(j)}}=s^{\gamma^{(j)}},\qquad\lim\limits_{r\to\infty}\frac{b_{\lfloor rs\rfloor}^{(j)}-b_{r}^{(j)}}{a_{r}^{(j)}}=\frac{s^{\gamma^{(j)}}-1}{\gamma^{(j)}},$ (2.1)
where the second limit is interpreted as $\log(s)$ if $\gamma^{(j)}=0$. Moreover, for $r\to\infty$,
$\bm{Z}_{r}=\frac{\bm{M}_{r}-\bm{b}_{r}}{\bm{a}_{r}}\rightsquigarrow\bm{Z}\sim G,$ (2.2)
where $G$ denotes a $d$-variate extreme-value distribution with marginal c.d.f.s $G_{\gamma^{(1)}},\dots,G_{\gamma^{(d)}}$ (with $G_{\gamma}(x)=\exp(-(1+\gamma x)^{-1/\gamma})\bm{1}(1+\gamma x>0)$ the cdf of the GEV($\gamma$)-distribution) and where division by the vector $\bm{a}_{r}$ is understood componentwise, that is, $\bm{Z}_{r}=(Z_{r}^{(1)},\dots,Z_{r}^{(d)})$ with $Z_{r}^{(j)}=\{\max(X_{1}^{(j)},\ldots,X_{r}^{(j)})-b_{r}^{(j)}\}/{a_{r}^{(j)}}$ for $j\in\{1,\dots,d\}$.
Note that (2.1) and (2.2) may for instance be deduced from Leadbetter’s $D(u_{n})$ condition, a domain-of-attraction condition on the associated i.i.d. sequence with stationary distribution equal to that of $\bm{X}_{0}$ and a weak requirement on the convergence of the c.d.f. of $\bm{Z}_{r},$ see Theorem 10.22 in Beirlant et al., (2004).
Throughout this paper, we assume to observe $\bm{X}_{1},\ldots,\bm{X}_{n}$, an excerpt from a $d$-dimensional time series $(\bm{X}_{t})_{t\in\mathbb{Z}}$ satisfying Condition 2.1. Block maxima will be formed with respect to a block size parameter $r=r_{n}$ converging to infinity such that $r=o(n)$. For establishing the subsequent limit results, the serial dependence of $(\bm{X}_{t})_{t}$ will additionally be controlled via the Rosenblatt mixing coefficient (Bradley, , 2005). For two sigma-fields $\mathcal{F}_{1},\mathcal{F}_{2}$ on a probability space $(\Omega,\mathcal{A},\mathbb{P})$, let $\alpha(\mathcal{F}_{1},\mathcal{F}_{2})=\sup_{A\in\mathcal{F}_{1},B\in\mathcal{F}_{2}}|\mathbb{P}(A\cap B)-\mathbb{P}(A)\mathbb{P}(B)|.$ For positive integer $p$, let $\alpha(p):=\alpha\big{(}\sigma((\bm{X}_{t})_{t\leq 0}),\sigma((\bm{X}_{t})_{t\geq p})\big{)}$, with $\sigma(\cdot)$ denoting the sigma-field generated by its argument.
Condition 2.2 (Block size and mixing).
For the block size sequence $(r_{n})_{n}$ it holds that, as $n\to\infty$:
  1. (a)
    $r_{n}\to\infty$ and $r_{n}=o(n).$
  2. (b)
    There exists a sequence $(\ell_{n})_{n}\subset\mathbb{N}$ such that $\ell_{n}\rightarrow\infty$, $\ell_{n}=o(r_{n})$, $\frac{r_{n}}{\ell_{n}}\alpha(\ell_{n})=o(1)$, and $\frac{n}{r_{n}}\alpha(\ell_{n})=o(1)$.
  3. (c)
    $\bigl{(}\frac{n}{r_{n}}\bigr{)}^{1+\omega}\alpha(r_{n})=o(1)$ for some $\omega>0$.
2.2 Empirical means of rescaled disjoint or sliding block maxima
A central theoretical ingredient for establishing weak limit results on estimators based on the block maxima method is the weak convergence of either the tail quantile process or of empirical means of (unobservable) rescaled block maxima. Respective results on the former can be found in Theorem 2.1 in Ferreira and de Haan, (2015), while the latter approach was taken in Theorems 2.6/B.1 in Bücher and Zanger, (2023), in the proof of Theorem 2.6 in Bücher and Segers, 2018a or in Theorem 2.4 in Zou et al., (2021). Throughout this paper, we also follow the latter approach, and for completeness, we summarize the essence of the respective results in a theorem and briefly summarize and discuss potential statistical applications.
[…]
$\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{mb}})}-P_{r}\big{)},\quad\mathbb{G}_{n,r}^{({\operatorname{mb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{mb}})}-P\big{)},$
where $\mathbb{P}_{n,r}^{({\operatorname{mb}})}=\frac{1}{n}\sum_{i=1}^{n}\delta_{\bm{Z}_{r,i}^{({\operatorname{mb}})}}$, $P_{r}=\mathbb{P}(\bm{Z}_{r}\in\cdot)$ and $P=\mathbb{P}(\bm{Z}\in\cdot)$. Finally, let
$G_{\xi}(\bm{x},\bm{y})=G(\bm{x})^{\xi\wedge 1}G(\bm{y})^{\xi\wedge 1}G(\bm{x}\wedge\bm{y})^{1-(\xi\wedge 1)},\qquad\bm{x},\bm{y}\in\mathbb{R}^{d},$ (2.3)
and note that $G_{\xi}$ is a $2d$-variate extreme-value distribution; see Formula (3.8) and its discussion in Bücher and Staud, (2024) for further details.
Theorem 2.3.
Under Conditions 2.1 and 2.2, for any finite set of real valued functions $h_{1},\dots,h_{q}$ satisfying the integrability Condition 10.1(a) with $\nu>2/\omega$ where $\omega$ is from Condition 2.2, we have, writing $\bm{h}=(h_{1},\dots,h_{q})^{\top}$,
$\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}\bm{h}=\big{(}\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}h_{1},\dots,\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}h_{q}\big{)}^{\top}\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma_{\bm{h}}^{({\operatorname{mb}})}),\qquad{\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\},$
where
$\big{(}\Sigma_{\bm{h}}^{({\operatorname{db}})}\big{)}_{j,j^{\prime}=1}^{q}=\operatorname{Cov}\big{(}h_{j}(\bm{Z}),h_{j^{\prime}}(\bm{Z})\big{)},\qquad\big{(}\Sigma_{\bm{h}}^{({\operatorname{sb}})}\big{)}_{j,j^{\prime}=1}^{q}=2\int_{0}^{1}\operatorname{Cov}\big{(}h_{j}(\bm{Z}_{1,\xi}),h_{j^{\prime}}(\bm{Z}_{2,\xi})\big{)}\,\mathrm{d}\xi,$ (2.4)
with $\bm{Z}\sim G$ from Condition 2.1 and $(\bm{Z}_{1,\xi},\bm{Z}_{2,\xi})\sim G_{\xi}$ from (2.3). Moreover, $\Sigma_{\bm{h}}^{({\operatorname{sb}})}\leq_{L}\Sigma_{\bm{h}}^{({\operatorname{db}})}$, where $\leq_{L}$ denotes the Loewner ordering. Finally, if the bias Condition 10.2 holds, then
$\mathbb{G}_{n,r}^{({\operatorname{mb}})}\bm{h}=\big{(}\mathbb{G}_{n,r}^{({\operatorname{mb}})}h_{1},\dots,\mathbb{G}_{n,r}^{({\operatorname{mb}})}h_{q}\big{)}^{\top}\rightsquigarrow\mathcal{N}_{q}(\bm{B}_{\bm{h}},\Sigma_{\bm{h}}^{({\operatorname{mb}})}).$
Proof.
The proof is a simplified version of our proof of Theorem 3.2 below, whence we omit further details. Under slightly different conditions, a proof can be found in Theorem B.1 in Bücher and Zanger, (2023) or Theorem 8.7 in Bücher and Staud, (2024). ∎
In statistical applications, one is typically interested in distributional approximations for the estimation error of general statistics depending on the observable block maxima $\bm{M}_{r,i}$ (here and in the following, we omit the upper index ‘${\operatorname{mb}}$’). As discussed in the following remark, such approximations can be deduced from Theorem 2.3.
Remark 2.4 (Normal approximations for statistics depending on observable block maxima).
Suppose that $\hat{\bm{\theta}}_{n}={\bm{\varphi}}_{n}(\bm{M}_{r,1},\dots,\bm{M}_{r,n})$ is some estimator of interest for a target parameter ${\bm{\theta}}_{r}\in\mathbb{R}^{p}$ (with $p\in\mathbb{N}$), where ${\bm{\varphi}}_{n}$ is a measurable function on $(\mathbb{R}^{d})^{n}$ with values in $\mathbb{R}^{p}$. As discussed in Example 2.5 below, for many functions ${\bm{\varphi}}_{n}$ of practical interest there exists a positive integer $q$ and functions ${\bm{\psi}}_{n}\colon(\mathbb{R}^{d})^{n}\to\mathbb{R}^{q}$, $A\colon(0,\infty)^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{p\times q}$ and $B\colon\mathbb{R}^{d}\times(0,\infty)^{d}\to\mathbb{R}^{p}$ such that
${\bm{\varphi}}_{n}(\bm{m}_{1},\dots,\bm{m}_{n})=A(\bm{a},\bm{b}){\bm{\psi}}_{n}\Big{(}\frac{\bm{m}_{1}-\bm{b}}{\bm{a}},\dots,\frac{\bm{m}_{n}-\bm{b}}{\bm{a}}\Big{)}+B(\bm{a},\bm{b})$ (2.5)
for all $\bm{m}_{1},\dots,\bm{m}_{n}\in\mathbb{R}^{d}$ and $\bm{b}\in\mathbb{R}^{d},\bm{a}\in(0,\infty)^{d}$. In such a case, if we further assume that there exists a sequence ${\bm{\vartheta}}_{r}\in\mathbb{R}^{q}$ solving the equations
${\bm{\theta}}_{r}=A(\bm{a}_{r},\bm{b}_{r}){\bm{\vartheta}}_{r}+B(\bm{a}_{r},\bm{b}_{r})$ (2.6)
(again, see below for examples), we immediately obtain that the estimation error of $\hat{\bm{\theta}}_{n}$ can be written as
$\hat{\bm{\theta}}_{n}-{\bm{\theta}}_{r}=A(\bm{a}_{r},\bm{b}_{r})(\hat{\bm{\vartheta}}_{n}-{\bm{\vartheta}}_{r}),$ (2.7)
where $\hat{\bm{\vartheta}}_{n}:={\bm{\psi}}_{n}(\bm{Z}_{r,1},\dots,\bm{Z}_{r,n})$ is a function of the rescaled block maxima. The difference $\hat{\bm{\vartheta}}_{n}-{\bm{\vartheta}}_{r}$ often allows for a linearization (for instance, by the delta-method): there exist real-valued functions $h_{1},\dots,h_{q}$ such that
$\sqrt{\frac{n}{r}}(\hat{\bm{\vartheta}}_{n}-{\bm{\vartheta}}_{r})=\bar{\mathbb{G}}_{n,r}(h_{1},\dots,h_{q})^{\top}+R_{n}$ (2.8)
for some remainder $R_{n}=o_{\mathbb{P}}(1)$. Loosely spoken, if the error term $R_{n}$ in the previous display is sufficiently small, the previous two displays and Theorem 2.3 imply the normal approximation
$\hat{\bm{\theta}}_{n}-{\bm{\theta}}_{r}=\sqrt{\frac{r}{n}}A(\bm{a}_{r},\bm{b}_{r})\big{\{}\bar{\mathbb{G}}_{n,r}(h_{1},\dots,h_{q})^{\top}+R_{n}\big{\}}\approx_{d}\mathcal{N}_{p}\Big{(}0,\frac{r}{n}A(\bm{a}_{r},\bm{b}_{r})\Sigma_{\bm{h}}A(\bm{a}_{r},\bm{b}_{r})^{\top}\Big{)},$ (2.9)
where $\Sigma_{\bm{h}}$ is the matrix from Theorem 2.3. More precisely, in the case where the matrix $A$ is invertible (in particular, $p=q$), we obtain the more precise result
$A(\bm{a}_{r},\bm{b}_{r})^{-1}\sqrt{\frac{n}{r}}(\hat{\bm{\theta}}_{r}-{\bm{\theta}}_{r})=\bar{\mathbb{G}}_{n,r}(h_{1},\dots,h_{q})^{\top}+o_{\mathbb{P}}(1)\rightsquigarrow\mathcal{N}_{p}(\bm{0},\Sigma_{\bm{h}}).$ (2.10)
Example 2.5.
(i) Empirical variance: consider the case $d=1$ and the empirical variance $\hat{\theta}_{n}=n^{-1}\sum_{i=1}^{n}(M_{r,i}-\bar{M}_{r,n})^{2},$ where $\bar{M}_{r,n}:=n^{-1}\sum_{i=1}^{n}M_{r,i}$, considered as an estimator for $\theta_{r}:=\operatorname{Var}(M_{r,1})$. In that case, (2.5) is met with $p=1,q=2$, ${\bm{\psi}}_{n}(z_{1},\dots,z_{n})=((\frac{1}{n}\sum_{i=1}^{n}z_{i})^{2},\frac{1}{n}\sum_{i=1}^{n}z_{i}^{2})^{\top}$, $B(a,b)=0$ and $A(a,b)=(-a^{2},a^{2})$. Moreover, ${\bm{\vartheta}}_{r}=(\operatorname{E}[Z_{r,1}]^{2},\operatorname{E}[Z_{r,1}^{2}])^{\top}$ is a solution to (2.6) satisfying (2.8): indeed, with $h_{1}(x)=x,h_{2}(x)=x^{2}$, and using $P_{r}h_{1}\to Ph_{1}$, we may write
$\sqrt{\frac{n}{r}}(\hat{\bm{\vartheta}}_{n}-{\bm{\vartheta}}_{r})=\sqrt{\frac{n}{r}}\begin{pmatrix}(\mathbb{P}_{n,r}h_{1})^{2}-(P_{r}h_{1})^{2}\\ \mathbb{P}_{n,r}h_{2}-P_{r}h_{2}\end{pmatrix}$ $=\sqrt{\frac{n}{r}}\begin{pmatrix}(2Ph_{1}+o_{\mathbb{P}}(1))(\mathbb{P}_{n,r}h_{1}-P_{r}h_{1})\\ \mathbb{P}_{n,r}h_{2}-P_{r}h_{2}\end{pmatrix}$
$=\bar{\mathbb{G}}_{n,r}((2Ph_{1})\cdot h_{1},h_{2})^{\top}+o_{\mathbb{P}}(1).$
Assembling terms and solving for $\hat{\theta}_{n}-\theta_{r}$, we obtain $\sqrt{\frac{n}{r}}a_{r}^{-2}(\hat{\theta}_{n}-\theta_{r})=\bar{\mathbb{G}}_{n,r}(h_{2}-(2Ph_{1})\cdot h_{1})+o_{\mathbb{P}}(1).$ The asymptotic variance of the limiting distribution has been explicitly calculated in Bücher and Staud, (2024), see their Corollary 4.1 and their Equations (C.1) and (C.2).
(ii) Probability weighted moments: consider the case $d=1$ and let ${\bm{\varphi}}_{n}:\mathbb{R}^{n}\to\mathbb{R}^{p}$ denote the vector containing the first $p$ empirical probability weighted moments. In that case, Equation (2.5) is met with $q=p$, ${\bm{\psi}}_{n}={\bm{\varphi}}_{n}$, $A(a,b)=\operatorname{diag}(a,\dots,a)$ and $B(a,b)=\operatorname{diag}(b,b/2,\dots,b/(p+1))$; see Formula (A.7) in Bücher and Zanger, (2023). Moreover, Equation (2.8) can be deduced from (A.5) in that paper (for $p=3$), with functions $h_{1}(x)=x$, $h_{2}(x)=xG_{\gamma}(x)+\int_{x}^{\infty}z\,\mathrm{d}G_{\gamma}(z)$ and $h_{3}(x)=xG_{\gamma}^{2}(x)+2\int_{x}^{\infty}zG_{\gamma}(z)\,\mathrm{d}G_{\gamma}(z)$. Finally, up to the treatment of a bias term, (2.10) corresponds to the assertion in their Theorem 3.2.
(iii) Pseudo maximum likelihood estimation in the heavy tailed case: the main steps from Remark 2.4 also apply in a slightly modified setting tailored to the heavy tailed case; details are worked out in Section 5 below.
In statistical applications, results like those in (2.9) are typically used as a starting point for inference, for instance in the form of confidence intervals for $\hat{\bm{\theta}}_{n}$. Routinely, such intervals would be based on normal approximations involving consistent estimators for the variances on the right-hand side of those displays, which in our case requires estimation of $\bm{a}_{r},\bm{b}_{r}$ and $\Sigma_{\bm{h}}=\Sigma_{\bm{h}}^{({\operatorname{mb}})}$. This can be a complicated task, especially for the sliding block maxima method in view of the complicated formula for the limiting variance. Alternatively, one can rely on bootstrap approximations instead of normal approximations, which can avoid the need for variance estimation. For the disjoint blocks case, the proof of Theorem 2.3 shows that the disjoint block maxima can essentially be considered independent (despite the serial dependence in the underlying observations). This suggests that the standard bootstrap based on resampling with replacement from the disjoint block maxima should be consistent; a conjecture that will be confirmed in Section 4. In the sliding block maxima case, however, the standard bootstrap cannot work since it destroys the serial dependence in the sliding block maxima sample. In classical time series analysis, a remedy consists of the block bootstrap, where blocks of successive observations are drawn with replacement. In our case, this heuristically requires the choice of a smoothing parameter sequence that is of larger order than the block size; see Remark 4.3 for details on the inconsistency of a naive approach to bootstrapping sliding block maxima. An alternative approach relies on the circular block maxima method as introduced in the next section.
3 The circular block maxima sample
In this section we introduce the circular block maxima method. We show that respective empirical means yield the same asymptotic variance as the sliding block maxima counterparts. The approach is interesting in its own (particularly for computational reasons), but most importantly for this paper it suggests a straightforward bootstrap approach for bootstrapping sliding block maxima that will be discussed in Section 4.
Formally, given a sample $\mathcal{X}_{n}=(\bm{X}_{1},\dots,\bm{X}_{n})$ as before, the circular block maxima sample
$\mathcal{M}_{n,r}^{({\operatorname{cb}})}=(\bm{M}_{r,1}^{({\operatorname{cb}})},\dots,\bm{M}_{r,n}^{({\operatorname{cb}})})$
is a new sample of size $n$ containing suitable block maxima that asymptotically follow the distribution $G$ in (2.2). The sample depends on two parameters: the block length parameter $r$ and an integer-valued parameter $k$ (for instance, $k=2$ or $k=3$) that determines the length of the interval over which we apply the circular maximum operation. Details for the univariate case are provided in Figure 1, where the sampling period $\{1,\dots,n\}$ has been decomposed into $m(k):=m_{n}(k):=n/(kr)$ blocks of size $kr$, say $I_{kr,1},\dots,I_{kr,m(k)}$ with $I_{kr,i}=\{(i-1)kr+1,\dots,ikr\}$, and where we assume that $m(k)\in\mathbb{N}$ for simplicity. Note that every observation $X_{s}$ appears within exactly $r$ maximum-operations, and hence has the same chance to become a block maximum (this is not the case for the plain sliding block maxima sample, where, for instance, the very first observation $X_{1}$ can appear only once as sliding block maximum). This observation is in fact the main motivation for the circularization approach within each $kr$-block.
[…]
Figure 1: Illustration of calculating the circular block maxima sample.
More formally, the notations used in Figure 1 are defined as follows: for a given vector $\bm{W}=(\bm{w}_{1},\ldots,\bm{w}_{n})\in(\mathbb{R}^{d})^{n}$ and $i\in\{1,\dots,m(k)\}$, we let
$\pi_{kr,i}(\bm{W})$ $=(\bm{w}_{s})_{s\in I_{kr,i}}=\big{(}\bm{w}_{(i-1)kr+1},\dots,\bm{w}_{ikr}\big{)}$
denote the projection of $\bm{W}$ to its coordinates defined by the $i$-th $kr$-block. For $\bm{V}=(\bm{v}_{1},\dots,\bm{v}_{q})\in(\mathbb{R}^{d})^{q}$ with $q\in\mathbb{N}_{\geq r}$, consider the sliding-maxima-operation
$\operatorname{slid{\text{-}}max}(\bm{V}\mid r)$ $=\Big{(}\max_{s\in[1:r]}\bm{v}_{s},\max_{s\in[2:r+1]}\bm{v}_{s},\dots,\max_{s\in[q-r+1:q]}\bm{v}_{s}\Big{)}\in(\mathbb{R}^{d})^{q-r+1},$
where maxima over vectors in $\mathbb{R}^{d}$ are to be understood componentwise and where $[i:j]=\{i,i+1,\dots,j\}$. Next, for $\bm{U}=(\bm{u}_{1},\dots,\bm{u}_{kr})\in(\mathbb{R}^{d})^{kr}$, the circularization function is defined as
$\operatorname{circ}(\bm{U}\mid r)$ $:=\big{(}\operatorname{circ}_{1}(\bm{U}\mid r),\dots,\operatorname{circ}_{kr+r-1}(\bm{U}\mid r)\big{)}:=(\bm{u}_{1},\dots,\bm{u}_{kr},\bm{u}_{1},\dots,\bm{u}_{r-1})\in(\mathbb{R}^{d})^{kr+r-1}.$
Finally, we define $\operatorname{circ{\text{-}}max}(\bm{U}\mid r):=\operatorname{slid{\text{-}}max}(\operatorname{circ}(\bm{U}\mid r)\mid r)$, that is,
$\operatorname{circ{\text{-}}max}(\bm{U}\mid r)$ $=\big{(}\operatorname{circ{\text{-}}max}_{1}(\bm{U}\mid r),\dots,\operatorname{circ{\text{-}}max}_{kr}(\bm{U}\mid r)\big{)}$
$=\Big{(}\max_{s\in[1:r]}\operatorname{circ}_{s}(\bm{U}\mid r),\max_{s\in[2:r+1]}\operatorname{circ}_{s}(\bm{U}\mid r),\dots,\max_{s\in[kr:kr+r-1]}\operatorname{circ}_{s}(\bm{U}\mid r)\Big{)}\in(\mathbb{R}^{d})^{kr}.$
It is constructive to take a closer look at the choices $k=1$ and $k=n/r$ (the number of disjoint blocks). In the former case, the circmax-sample is the same as the disjoint block maxima sample, but with every observation repeated exactly $r$ times. In the latter case, the first $n-r+1$ observations of the circmax-sample coincide with the sliding block maxima sample. New samples are obtained for every other choice of $k$, and in the subsequent developments we will mostly be interested in choices like $k=2$ or $k=3$.
A key feature of the circmax-sample consists of the fact that it can be stored and evaluated efficiently. More precisely, for each $kr$ block, the number of distinct circmax-values in that block is typically very small (typically around 10 for realistic block and sample sizes as observed in the simulation study); for instance, the largest observation in each block is necessarily appearing exactly $r$ times. As such, the entire circmax-sample can also be regarded as a weighted sample whose (random) size corresponds to the number of distinct values in the circmax-sample.
The following central result shows that the circmax-sample can be considered as a sample from the ‘correct’-limit distribution $G$ from (2.2) (a similar result on joint convergence of two circmax-observations can be found in Proposition 11.1). For $s\in\{1,\dots,n\}$, define
$\bm{Z}_{r,s}^{({\operatorname{cb}})}:=\frac{\bm{M}_{r,s}^{({\operatorname{cb}})}-\bm{b}_{r}}{\bm{a}_{r}}.$
Note that we suppress the dependence on $k$ for notational convenience.
Proposition 3.1 (Weak convergence of circular block maxima).
Suppose that Condition 2.1 and 2.2(a) are met and that $\alpha(r_{n})\to 0$. Then, for every fixed $k\in\mathbb{N}$ and $\xi\in[0,k)$, we have
$\bm{Z}_{r,1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\rightsquigarrow\bm{Z}\sim G\qquad(n\to\infty).$
Due to Proposition 3.1, statistical methods based on the circular block maxima method may be expected to work, asymptotically. As discussed in Section 2.2, a central ingredient for studying respective methods is weak convergence of empirical means. Adopting the notation from that section, we denote the empirical processes associated with the normalized sample $\bm{Z}_{r,1}^{({\operatorname{cb}})},\dots,\bm{Z}_{r,n}^{({\operatorname{cb}})}$ by
$\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{cb}})}-P_{n,r}^{({\operatorname{cb}})}\big{)},\quad\tilde{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{cb}})}-P_{r}\big{)},\quad\mathbb{G}_{n,r}^{({\operatorname{cb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{cb}})}-P\big{)},$ (3.1)
where $\mathbb{P}_{n,r}^{({\operatorname{cb}})}=\frac{1}{n}\sum_{s=1}^{n}\delta_{\bm{Z}_{r,s}^{({\operatorname{cb}})}}$, $P_{n,r}^{({\operatorname{cb}})}=\frac{1}{kr}\sum_{s=1}^{kr}\mathbb{P}(\bm{Z}_{r,s}^{({\operatorname{cb}})}\in\cdot)$, $P_{r}=\mathbb{P}(\bm{Z}_{r,1}\in\cdot)$ and $P=\mathbb{P}(\bm{Z}\in\cdot)$. The following result can be regarded as a circmax-counterpart of Theorem 2.3.
Theorem 3.2.
Suppose that Conditions 2.1 and 2.2 are met. Then, for fixed $k\in\mathbb{N}_{\geq 2}$ and any finite set of real valued functions $h_{1},\dots,h_{q}$ satisfying the integrability Condition 10.1(b) with $\nu>2/\omega$ with $\omega$ from Condition 2.2, we have
$\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}\bm{h}\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma_{\bm{h}}^{({\operatorname{sb}})})$
and $\lim_{n\to\infty}\operatorname{Cov}(\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}\bm{h})=\Sigma_{\bm{h}}^{({\operatorname{sb}})}$ with $\Sigma_{\bm{h}}^{({\operatorname{sb}})}$ from (2.4). Moreover, if the bias Condition 10.3 holds, then
$\tilde{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}\bm{h}\rightsquigarrow\mathcal{N}_{q}(\tfrac{D_{\bm{h},k}+E_{\bm{h}}}{k},\Sigma_{\bm{h}}^{({\operatorname{sb}})})$
(with $D_{\bm{h},k}$ often equal to zero, see Remark 3.3), and if additionally the bias Condition 10.2 is met, then
$\mathbb{G}_{n,r}^{({\operatorname{cb}})}\bm{h}\rightsquigarrow\mathcal{N}_{q}(B_{\bm{h}}+\tfrac{D_{\bm{h},k}+E_{\bm{h}}}{k},\Sigma_{\bm{h}}^{({\operatorname{sb}})}).$
Remark 3.3.
The bias term $D_{\bm{h},k}+E_{\bm{h}}$ is due to the fact that the rescaled circular block maxima $\max(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr},\bm{X}_{1},\ldots,\bm{X}_{t})$ with index $s=(k-1)r+1+t$ for some $t\in\{1,\dots,r-1\}$ are not equal in distribution to a generic block maximum variable $\bm{M}_{r}$. This bias has two sources: $D_{\bm{h},k}$ is a measure for the average (over $t$) difference induced by the approximation of $\max(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr},\bm{X}_{1},\ldots\bm{X}_{t})$ by $\max(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr},\tilde{\bm{X}}_{1},\ldots,\tilde{\bm{X}}_{t})$ (with $(\tilde{\bm{X}}_{t})_{t}$ an independent copy of $(\tilde{\bm{X}}_{t})_{t}$), while $E_{\bm{h}}$ is a measure for the average difference induced by the approximation of $\max(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr},\tilde{\bm{X}}_{1},\ldots,\tilde{\bm{X}}_{t})$ by that of a generic block maximum variable $\bm{M}_{r}$. In Lemma 10.4 we provide simple sufficient conditions for beta mixing time series that imply $D_{\bm{h},k}=B_{\bm{h}}=0$.
The proof of Theorem 3.2 (see in particular Proposition 11.1) shows that the sample of $kr$-blocks containing the circular block maxima can be considered asymptotically independent. This is akin to the sample of disjoint block maxima and, as discussed after Theorem 2.3, it suggests to define bootstrap versions of $\mathbb{G}_{n,r}^{({\operatorname{cb}})}$ by drawing $m(k)$-times from the sample of $kr$-blocks with replacement (which is computationally efficient: recall that all circular block maxima in a $kr$-block could be efficiently stored as a weighted sample with typically very few observations).
4 Bootstrapping block maxima estimators
Throughout this section, we discuss bootstrap approximations for the empirical processes from Section 3. For completeness, we also cover the disjoint block maxima case, as it can be identified with the circular block maxima sample for $k=1$.
For fixed $k\in\mathbb{N}$, consider the circmax-sample $\mathcal{M}_{n,r}^{({\operatorname{cb}})}$. Independent of the observations, let $\bm{W}_{m(k)}=(W_{m(k),1},\dots,W_{m(k),m(k)})=(W_{1},\dots,W_{m(k)})$ be multinomially distributed with $m(k)$ trials and class probabilities $(m(k)^{-1},\dots,m(k)^{-1})$. Given the sample $\mathcal{M}_{n,r}^{({\operatorname{cb}})}$, the bootstrap sample $\mathcal{M}_{n,r}^{({\operatorname{cb}}),*}$ is obtained by repeating the observations $(M_{r,s}^{({\operatorname{cb}})})_{s\in I_{kr,i}}$ from the $i$th $kr$-block exactly $W_{m(k),i}$-times, for every $i=1,\dots,m(k)$. The respective empirical measure of the rescaled bootstrap sample $\bm{Z}_{r,1}^{({\operatorname{cb}}),*},\dots,\bm{Z}_{r,n}^{({\operatorname{cb}}),*}$ with $\bm{Z}_{r,1}^{({\operatorname{cb}}),*}=(\bm{M}_{r,s}^{({\operatorname{cb}}),*}-\bm{b}_{r})/\bm{a}_{r}$ can then be written as
^$\hat{\mathbb{P}}_{n,r}^{({\operatorname{cb}}),*}=\frac{1}{n}\sum_{s=1}^{n}\delta_{\bm{Z}_{r,s}^{({\operatorname{cb}}),*}}=\frac{1}{n}\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\delta_{\bm{Z}_{r,s}^{({\operatorname{cb}})}},$
and the associated (centered) empirical process is given by
$\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}=\sqrt{\frac{n}{r}}(\hat{\mathbb{P}}_{n,r}^{({\operatorname{cb}}),*}-\mathbb{P}_{n,r}^{({\operatorname{cb}})}).$ (4.1)
Note that the conditional distribution of $\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}$ given the data $\bm{X}_{1},\dots,\bm{X}_{n}$ can in practice be approximated to an arbitrary precision based on repeated sampling of $\bm{W}_{m(k)}$ (if $\bm{a}_{r},\bm{b}_{r}$ were known; we discuss extensions below). The following result shows that the bootstrap process $\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}$ provides a consistent distributional approximation for $\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}$ and for $\bar{\mathbb{G}}_{n,r}^{({\operatorname{sb}})}$.
Theorem 4.1 (Asymptotic validity of the circmax-resampling bootstrap).
Suppose Conditions 2.1 and 2.2 are met. Then, for fixed $k\in\mathbb{N}_{\geq 2}$ and $\bm{h}$ satisfying the integrability Condition 10.1(b) with $\nu>2/\omega$ with $\omega$ from Condition 2.2, we have, for ${\operatorname{mb}}\in\{{\operatorname{sb}},{\operatorname{cb}}\}$,
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}\bm{h}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}\bm{h}\big{)}\Big{)}=o_{\mathbb{P}}(1)$
as $n\to\infty$, where $d_{K}$ denotes the Kolmogorov metric for probability measures on $\mathbb{R}^{q}$. As a consequence, if the bias Condition 10.2 is met and if $r$ is chosen sufficiently large to guarantee that $B_{\bm{h}}=0$, we have
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}\bm{h}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\mathbb{G}_{n,r}^{({\operatorname{sb}})}\bm{h}\big{)}\Big{)}=o_{\mathbb{P}}(1),$
and if also the bias Condition 10.3 is met with $E_{\bm{h}}=D_{\bm{h},k}=0$, we have
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}\bm{h}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\mathbb{G}_{n,r}^{({\operatorname{cb}})}\bm{h}\big{)}\Big{)}=o_{\mathbb{P}}(1).$
Remarkably, the consistency of the circmax bootstrap for the sliding block maxima method does not require the bias Condition 10.3 (nor $E_{\bm{h}}=D_{\bm{h},k}=0$ in case it is met).
Remark 4.2 (Bootstrapping the disjoint block maxima empirical process).
Consider the disjoint block maxima sample, whose empirical means coincide with empirical means based on the circular block maxima sample with $k=1$. The results from Theorem 4.1 continue to hold for $k=1$ ($D_{\bm{h},k}=E_{\bm{h}}=0$ is then immediate), but with ‘${\operatorname{sb}}$’ replaced by ‘${\operatorname{db}}$’ at all instances.
Remark 4.3 (Inconsistency of naive resampling of sliding block maxima).
It seems natural to bootstrap sliding block maxima empirical means by the sliding block maxima analogue of (4.1), that is, by
$\hat{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}=\sqrt{\frac{n}{r}}(\hat{\mathbb{P}}_{n,r}^{({\operatorname{sb}}),*}-P_{r}),\qquad\hat{\mathbb{P}}_{n,r}^{({\operatorname{sb}}),*}=\frac{1}{n}\sum_{s=1}^{n}\delta_{\bm{Z}_{r,s}^{({\operatorname{sb}}),*}}=\frac{1}{n}\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\delta_{\bm{Z}_{r,s}^{({\operatorname{sb}})}}.$
Note that $\hat{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}$ depends on $k$, which is suppressed from the notation. However, unlike $\mathbb{G}_{n,r}^{({\operatorname{cb}}),*}$, this process is inconsistent for both $\bar{\mathbb{G}}_{n,r}^{({\operatorname{sb}})}$ and $\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}$. Indeed, considering the case $d=q=1$ for simplicity, we prove in the appendix that
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}h\mid\mathcal{X}_{n}\big{)},\mathcal{N}(\bm{0},\Sigma_{h}^{(k)})\Big{)}=o_{\mathbb{P}}(1),$ (4.2)
where $\Sigma_{h}^{(k)}=\Sigma_{h}^{({\operatorname{sb}})}-\frac{2}{k}\int_{0}^{1}\xi\operatorname{Cov}(h(\bm{Z}_{1,\xi}),h(\bm{Z}_{2,\xi}))\,\mathrm{d}\xi.$ If $\operatorname{Var}(h(\bm{Z}))>0$, we have $\Sigma_{h}^{(k)}<\Sigma_{h}^{({\operatorname{sb}})}$ for any $k\in\mathbb{N}$ by Lemma 11.4, so in view of Theorem 3.2 the bootstrap process has a smaller asymptotic variance than needed for approximating $\bar{\mathbb{G}}_{n,r}^{({\operatorname{sb}})}$ or $\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}$. This circumstance is one of the main motivations for working with circular block maxima.
In statistical applications, as discussed after Theorem 2.3, we do not want to bootstrap the estimation error of empirical means involving the (unobservable) $\bm{Z}_{r,i}$, but the estimation error of general statistics depending on the block maxima $\bm{M}_{r,i}$ itself. We follow the general setting of Remark 2.4, that is, $\hat{\bm{\theta}}_{n}={\bm{\varphi}}_{n}(\bm{M}_{r,1},\dots,\bm{M}_{r,n})$ is some estimator of interest for a target parameter ${\bm{\theta}}_{r}\in\mathbb{R}^{p}$. If similar equations/linearizations as in (2.7) and (2.8) can be shown to hold for the (unobservable) circmax-bootstrap counterpart $\hat{\bm{\vartheta}}_{n}^{({\operatorname{cb}}),*}={\bm{\psi}}_{n}(\bm{Z}_{1,r}^{({\operatorname{cb}}),*},\dots,\bm{Z}_{n,r}^{({\operatorname{cb}}),*})$, that is, if
$\hat{\bm{\theta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\theta}}_{n}^{({\operatorname{cb}})}$ $=A(\bm{a}_{r},\bm{b}_{r})(\hat{\bm{\vartheta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\vartheta}}_{n}^{({\operatorname{cb}})}),$ (4.3)
$\sqrt{\frac{n}{r}}(\hat{\bm{\vartheta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\vartheta}}_{n}^{({\operatorname{cb}})})$ $=\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}(h_{1},\dots,h_{q})^{\top}+o_{\mathbb{P}}(1)$ (4.4)
then, loosely speaking, the distributional approximation $\mathbb{G}_{n,r}^{({\operatorname{cb}})}\bm{h}\approx_{d}(\mathbb{G}_{n,r}^{({\operatorname{cb}}),*}\bm{h}\mid\mathcal{X}_{n})$ from Theorem 4.1 carries over to the distributional approximation $\hat{\bm{\theta}}_{n}^{({\operatorname{cb}})}-{\bm{\theta}}_{r}\approx_{d}(\hat{\bm{\theta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\theta}}_{n}^{({\operatorname{cb}})}\mid\mathcal{X}_{n})$. Formally, we have the following result.
Proposition 4.4.
Fix $k\in\mathbb{N}_{\geq 2}$ and assume that Conditions 2.1 and 2.2 are met. Suppose $\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}={\bm{\varphi}}_{n}(\bm{M}_{r,1}^{({\operatorname{mb}})},\dots,\bm{M}_{r,n}^{({\operatorname{mb}})})$ is some estimator of interest for a target parameter ${\bm{\theta}}_{r}\in\mathbb{R}^{p}$ such that (2.5), (2.6) and (2.8) (with $\bar{\mathbb{G}}_{n,r}=\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}$) is met for some function $\bm{h}=(h_{1},\dots,h_{q})^{\top}$ satisfying the conditions of Theorem 3.2. Moreover, assume that (4.3) and (4.4) is met. Then, if $\Sigma_{\bm{h}}^{({\operatorname{sb}})}$ is invertible, for ${\operatorname{mb}}\in\{{\operatorname{sb}},{\operatorname{cb}}\}$,
$d_{K}\Big{(}\mathcal{L}\big{(}\hat{\bm{\theta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\theta}}_{n}^{({\operatorname{cb}})}\mid\mathcal{X}_{n}\big{)},\mathcal{L}\big{(}\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}-{\bm{\theta}}_{r}\big{)}\Big{)}=o_{\mathbb{P}}(1).$
Proof.
By Equations (2.5), (2.6), (2.8) and (4.3), (4.4) we have
$\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}-{\bm{\theta}}_{r}$ $=A(\bm{a}_{r},\bm{b}_{r})\sqrt{\frac{r}{n}}\big{(}\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}\bm{h}+o_{\mathbb{P}}(1)\big{)}=:\sqrt{\frac{r}{n}}A(\bm{a}_{r},\bm{b}_{r})S_{n},$
$\hat{\bm{\theta}}_{n}^{({\operatorname{cb}}),*}-\hat{\bm{\theta}}_{n}^{({\operatorname{cb}})}$ $=A(\bm{a}_{r},\bm{b}_{r})\sqrt{\frac{r}{n}}\big{(}\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}},\ast)}\bm{h}+o_{\mathbb{P}}(1)\big{)}=:\sqrt{\frac{r}{n}}A(\bm{a}_{r},\bm{b}_{r})S_{n}^{*}.$
By Theorems 2.3 and 3.2 we have $S_{n}\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma_{\bm{h}}^{({\operatorname{sb}})})$. Theorem 4.1 implies $d_{K}(\mathcal{L}(S_{n}^{*}\mid\mathcal{X}_{n}),\mathcal{L}(S_{n}))=o_{\mathbb{P}}(1).$ The assertion then follows from Lemma 11.9. ∎
The results from Proposition 4.4 are sufficient for showing that basic bootstrap confidence intervals (Davison and Hinkley, , 1997) asymptotically hold their intended level. As a proof of concept, details on a specific example are given in the next section, see Corollary 5.6.
5 Application: bootstrapping the pseudo-maximum likelihood estimator for the Fréchet distribution
In this section, we provide details how the previous methods and results can be used for a specific important estimation problem. More precisely, we consider the univariate heavy tailed case, and restrict attention to the maximum likelihood estimator for parameters associated with a suitable version of the max-domain of attraction Condition 2.1.
Condition 5.1 (Fréchet Max-Domain of Attraction).
Let $(X_{t})_{t\in\mathbb{Z}}$ denote a strictly stationary univariate time series with continuous margins. There exists some $\alpha_{0}>0$ and some sequence $(\sigma_{r})_{r\in\mathbb{N}}\subset(0,\infty)$ such that
$\lim_{r\to\infty}\frac{\sigma_{\lfloor rs\rfloor}}{\sigma_{r}}=s^{1/\alpha_{0}},\,\,(s>0)\quad\text{ and }\quad\frac{\max(X_{1},\ldots,X_{r})}{\sigma_{r}}\rightsquigarrow Z\sim P_{\alpha_{0},1}\qquad(r\to\infty),$
where $P_{\alpha,\sigma}$ denotes the Fréchet-scale family on $(0,\infty)$ defined by its cdf $F_{\alpha,\sigma}(x)=\exp(-(x/\sigma)^{-\alpha})$; here $(\alpha,\sigma)\in(0,\infty)^{2}$.
Suppose $X_{1},\dots,X_{n}$ is an observed sample from $(X_{t})_{t\in\mathbb{Z}}$ as in Condition 5.1, and let $r=r_{n}$ denote a block length parameter. In view of the heuristics and results from Sections 2 and 3, the associated block maxima samples $\mathcal{M}_{n,r}^{({\operatorname{mb}})}$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}\}$ (with $k\in\mathbb{N}$ fixed) can all be considered approximate samples from $P_{\alpha_{0},\sigma_{r}}$. As in Bücher and Segers, 2018b ; Bücher and Segers, 2018a , this suggests to estimate $(\alpha_{0},\sigma_{r})$ by maximizing the independence Fréchet-log-likelihood, that is, we define
$\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}:=(\hat{\alpha}_{n}^{({\operatorname{mb}})},\hat{\sigma}_{n}^{({\operatorname{mb}})})^{\top}:=\operatornamewithlimits{\arg\max}_{{\bm{\theta}}=(\alpha,\sigma)\in(0,\infty)^{2}}\sum_{M_{i}\in\mathcal{M}_{n,r}^{({\operatorname{mb}})}}\ell_{\bm{\theta}}(M_{i}\vee c),$ (5.1)
where, for ${\bm{\theta}}=(\alpha,\sigma)^{\top}\in(0,\infty)^{2}$,
$\ell_{\bm{\theta}}(x)=\log(\alpha/\sigma)-(x/\sigma)^{-\alpha}-(\alpha+1)\log(x/\sigma),\quad x>0,$ (5.2)
denotes the log density of the Fréchet distribution $P_{(\alpha,\sigma)}$ and where $c>0$ denotes an arbitrary truncation constant. For the case ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\}$, the rescaled estimation error of $\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}$ is known to be asymptotically normal (and independent of $c$), that is, under suitable additional assumptions,
$\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{({\operatorname{mb}})}-\alpha_{0}\\ \hat{\sigma}_{n}^{({\operatorname{mb}})}/\sigma_{r_{n}}-1\end{pmatrix}\rightsquigarrow\mathcal{N}_{2}(\bm{\mu},\Sigma^{({\operatorname{mb}})}),$ (5.3)
for some $\bm{\mu}\in\mathbb{R}^{2}$ and some $\Sigma^{({\operatorname{mb}})}\in\mathbb{R}^{2\times 2}$ positive definite, see Bücher and Segers, 2018b ; Bücher and Segers, 2018a . Subsequently, we will extend these results to ${\operatorname{mb}}={\operatorname{cb}}$, and we will derive a consistent bootstrap scheme for the estimation error.
For fixed $k\in\mathbb{N}$, consider the circmax-sample $\mathcal{M}_{n,r}^{({\operatorname{cb}})}$. Independent of the observations, let $\bm{W}_{m(k)}=(W_{m(k),1},\dots,W_{m(k),m(k)})=(W_{1},\dots,W_{m(k)})$ be multinomially distributed with $m(k)$ trials and class probabilities $(m(k)^{-1},\dots,m(k)^{-1})$. As in Section 4, conditional on $\mathcal{M}_{n,r}^{({\operatorname{cb}})}$, the bootstrap sample $\mathcal{M}_{n,r}^{({\operatorname{cb}}),*}$ is obtained by repeating the observations $(M_{r,s}^{({\operatorname{cb}})})_{s\in I_{kr,i}}$ from the $i$th $kr$-block exactly $W_{m(k),i}$-times, for every $i=1,\dots,m(k)$. We are going to show that the independence Fréchet-log-likelihood
${\bm{\theta}}\mapsto\sum_{s=1}^{n}\ell_{\bm{\theta}}(M_{r,s}^{({\operatorname{cb}}),*}\vee c)=\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\ell_{\bm{\theta}}(M_{r,s}^{({\operatorname{cb}})}\vee c).$
has a unique-maximimizer (with probability converging to one), say $(\hat{\alpha}_{n}^{({\operatorname{cb}}),*},\hat{\sigma}_{n}^{({\operatorname{cb}}),*})$, and that the conditional distribution of the rescaled bootstrap-estimation error, $\sqrt{n/r}(\hat{\alpha}_{n}^{({\operatorname{cb}}),*}-\hat{\alpha}_{n}^{({\operatorname{cb}})},\hat{\sigma}_{n}^{({\operatorname{cb}}),*}/\hat{\sigma}_{n}-1)$, given the observations is close to the distribution of $\sqrt{n/r}(\hat{\alpha}_{n}^{({\operatorname{mb}})}-\alpha_{0},\hat{\sigma}_{n}^{({\operatorname{mb}})}/\sigma_{r_{n}}-1)$ for both ${\operatorname{mb}}={\operatorname{sb}}$ and ${\operatorname{mb}}={\operatorname{cb}}$. We also show how this result can be used to derive valid asymptotic confidence intervals.
A couple of conditions akin to those imposed in Bücher and Segers, 2018b ; Bücher and Segers, 2018a will be needed.
Condition 5.2 (All disjoint block maxima of size $\lfloor r_{n}/2\rfloor$ diverge).
For all $c>0$, the probability of the event that all disjoint block maxima of block size $\tilde{r}_{n}=\lfloor r_{n}/2\rfloor$ are larger than $c$ converges to one.
This is Condition 2.2 in Bücher and Segers, 2018a , and guarantees that the probability of the event that all (blocksize $r_{n}$) block maxima in $\mathcal{M}_{n,r}^{({\operatorname{mb}})}$ are larger than $c$ converges to one as well, for ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}\}$. This will guarantee that the Fréchet-log-likelihoods under consideration are well-defined, with probability converging to one. The following condition is Condition 3.4 in Bücher and Segers, 2018b ; recall that $M_{r,1}=\max(X_{1},\dots,X_{r})$.
Condition 5.3 (Moments).
There exists $\nu>2/\omega$ with $\omega$ from Condition 2.2 such that
$\limsup_{r\to\infty}\operatorname{E}[g_{\nu,\alpha_{0}}\big{(}(M_{r,1}\vee 1)/\sigma_{r}\big{)}]<\infty,$
where $g_{\nu,\alpha_{0}}(x):=\{x^{-\alpha_{0}}\bm{1}\{x\leq e\}+\log(x)\bm{1}\{x>e\}\}^{2+\nu}$.
Condition 5.4 (Bias).
There exists $c_{0}>0$ such that, for $j\in\{1,2,3\}$, the limits $C(f_{j}):=\lim_{n\to\infty}C_{n}(f_{j})$ and $C_{k}(f_{j}):=\lim_{n\to\infty}C_{n,k}(f_{j})$ exist, where
$C_{n}(f_{j})$ $:=\sqrt{\frac{n}{r}}\left(\operatorname{E}\big{[}f_{j}\big{(}(M_{r,1}\vee c_{0})/\sigma_{r}\big{)}\big{]}-P_{\alpha_{0},1}f_{j}\right)$
$C_{n,k}(f_{j})$ $:=\sqrt{\frac{n}{r}}\Big{(}\frac{1}{r}\sum_{s=(k-1)r+1}^{kr}\big{\{}\operatorname{E}\big{[}f_{j}\big{(}(M_{r,s}\vee c_{0})/\sigma_{r}\big{)}\big{]}-\operatorname{E}\big{[}f_{j}\big{(}(M_{r,1}\vee c_{0})/\sigma_{r}\big{)}\big{]}\big{\}}\Big{)}$
where $f_{j}:(0,\infty)\to\mathbb{R}$ are defined as
$f_{1}(x)=x^{-\alpha_{0}},\quad f_{2}(x)=x^{-\alpha_{0}}\log x,\quad f_{3}(x)=\log x.$ (5.4)
The first part of this condition on $C_{n}$ is Condition 3.5 in Bücher and Segers, 2018b , and provides control of the bias for those block maxima which are calculated based on $r$ successive observations. For the circmax-sample, the last $r$ observations in a $kr$-block of that sample are not of that form; their contribution to the bias is controlled by the condition on $C_{n,k}$.
Subsequently, we fix a truncation constant $c>0$, and we write
$\mathbb{G}_{n,r}^{({\operatorname{cb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{cb}})}-P_{(\alpha_{0},1)}\big{)},\quad\hat{\mathbb{G}}_{n}^{({\operatorname{cb}}),*}=\sqrt{\frac{n}{r}}(\hat{\mathbb{P}}_{n,r}^{({\operatorname{cb}}),\ast}-\mathbb{P}_{n,r}^{({\operatorname{cb}})})$
where
$\mathbb{P}_{n,r}^{({\operatorname{cb}})}=\frac{1}{n}\sum_{s=1}^{n}\delta_{(M_{r,s}^{({\operatorname{cb}})}\vee c)/\sigma_{r}},\qquad\hat{\mathbb{P}}_{n,r}^{({\operatorname{cb}}),*}=\frac{1}{n}\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\delta_{(M_{r,s}^{({\operatorname{cb}})}\vee c)/\sigma_{r}};$
the double use of notation with (3.1) and (4.1) should not cause any confusion. Furthermore, we suppress the dependence on $c$ in the notation, which is motivated by the fact that the limiting distribution does not depend on $c$, as shown in the following theorem.
Theorem 5.5.
Fix $k\in\mathbb{N}_{\geq 2}$ and $c>0$, and suppose that Conditions 2.2 and 5.15.4 are satisfied. Then, with probability tending to one, there exists a unique maximizer $(\hat{\alpha}_{n}^{({\operatorname{cb}})},\hat{\sigma}_{n}^{({\operatorname{cb}})})$ of the Fréchet log-likelihood ${\bm{\theta}}\mapsto\sum_{i=1}^{n}\ell_{\bm{\theta}}(M_{r,s}^{({\operatorname{cb}})}\vee c)$, and this maximizer satisfies
$\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{({\operatorname{cb}})}-\alpha_{0}\\ \hat{\sigma}_{n}^{({\operatorname{cb}})}/\sigma_{r_{n}}-1\end{pmatrix}=M(\alpha_{0})\mathbb{G}_{n,r}^{({\operatorname{cb}})}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)\rightsquigarrow\mathcal{N}_{2}(M(\alpha_{0})(\bm{C}+k^{-1}\bm{C}_{k}),\Sigma^{({\operatorname{sb}})}),$
where $\bm{C}=(C(f_{1}),C(f_{2}),C(f_{3}))^{\top},\bm{C}_{k}=(C_{k}(f_{1}),C_{k}(f_{2}),C_{k}(f_{3}))^{\top}$ and
$M(\alpha_{0})=\frac{6}{\pi^{2}}\begin{pmatrix}\alpha_{0}^{2}&\alpha_{0}(1-\gamma)&-\alpha_{0}^{2}\\ \gamma-1&-(\Gamma^{\prime\prime}(2)+1)/\alpha_{0}&1-\gamma\end{pmatrix},\quad\Sigma^{({\operatorname{sb}})}=\begin{pmatrix}0.4946\alpha_{0}^{2}&-0.3236\\ -0.3236&0.9578\alpha_{0}^{-2}\end{pmatrix}$
with $\Gamma(z)=\int_{0}^{\infty}t^{z-1}{e}^{-t}\,\mathrm{d}t$ the Euler Gamma function and $\gamma=0.5772\ldots$ the Euler-Mascheroni constant.
Moreover, also with probability tending to one, there exists a unique maximizer $(\hat{\alpha}_{n}^{({\operatorname{cb}}),*},\hat{\sigma}_{n}^{({\operatorname{cb}}),*})$ of ${\bm{\theta}}\mapsto\sum_{s=1}^{n}\ell_{\bm{\theta}}(M_{r,s}^{({\operatorname{cb}}),*}\vee c)$, and this maximizer satisfies
$\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{({\operatorname{cb}}),*}-\hat{\alpha}_{n}^{({\operatorname{cb}})}\\ \hat{\sigma}_{n}^{({\operatorname{cb}}),*}/\hat{\sigma}_{n}^{({\operatorname{cb}})}-1\end{pmatrix}=M(\alpha_{0})\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)\rightsquigarrow\mathcal{N}_{2}(0,\Sigma^{({\operatorname{sb}})}).$
Fix ${\operatorname{mb}}\in\{{\operatorname{sb}},{\operatorname{cb}}\}$. If $r_{n}$ is chosen sufficiently large such that, for $j\in\{1,2,3\}$, $C(f_{j})=0$ (${\operatorname{mb}}={\operatorname{sb}}$) or $C(f_{j})=C_{k}(f_{j})=0$ (${\operatorname{mb}}={\operatorname{cb}}$) in Condition 5.4, we have bootstrap consistency in the following sense:
$d_{K}\bigg{[}\mathcal{L}\bigg{(}\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{({\operatorname{cb}}),*}-\hat{\alpha}_{n}^{({\operatorname{cb}})}\\ \hat{\sigma}_{n}^{({\operatorname{cb}}),*}/\hat{\sigma}_{n}^{({\operatorname{cb}})}-1\end{pmatrix}\,\Big{|}\,\mathcal{X}_{n}\bigg{)},\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{({\operatorname{mb}})}-\alpha_{0}\\ \hat{\sigma}_{n}^{({\operatorname{mb}})}/\sigma_{r_{n}}-1\end{pmatrix}\bigg{]}=o_{\mathbb{P}}(1).$
The result in Theorem 5.5 allows for statistical inference on the parameters $\alpha_{0},\sigma_{r}$, for instance in the form of confidence intervals. We provide details on $\sigma_{r}$, using the circular block bootstrap approximation to the sliding block estimator: for $\beta\in(0,1)$, let $q_{\hat{\sigma}_{n}^{*}}(\beta)$ denote the $\beta$-quantile of the conditional distribution of $\hat{\sigma}_{n}^{({\operatorname{cb}}),*}$ given the data, that is, $q_{\hat{\sigma}_{n}^{*}}(\beta)=(F_{\hat{\sigma}_{n}^{*}})^{-1}(\beta)$, where $F_{\hat{\sigma}_{n}^{*}}(x)=\mathbb{P}(\hat{\sigma}_{n}^{({\operatorname{cb}}),*}\leq x\mid\mathcal{X}_{n})$ for $x\in\mathbb{R}$. Note that the quantile may be approximated to an arbitrary precision by repeated bootstrap sampling. Consider the basic bootstrap confidence interval (Davison and Hinkley, , 1997)
$I_{n,\sigma}^{({\operatorname{sb}},{\operatorname{cb}})}(1-\beta)$ $=\big{[}2\hat{\sigma}_{n}^{({\operatorname{sb}})}-q_{\hat{\sigma}_{n}^{*}}(1-\tfrac{\beta}{2}),2\hat{\sigma}_{n}^{({\operatorname{sb}})}-q_{\hat{\sigma}_{n}^{*}}(\tfrac{\beta}{2})\big{]}.$
Corollary 5.6.
Under the conditions of Theorem 5.5 with $r_{n}$ chosen sufficiently large such that, for $j\in\{1,2,3\}$, $C(f_{j})=0$ in Condition 5.4, we have, for any $\beta\in(0,1)$,
$\lim_{n\to\infty}\mathbb{P}(\sigma_{r_{n}}\in I_{n,\sigma}^{({\operatorname{sb}},{\operatorname{cb}})}(1-\beta))=1-\beta.$
The above result only concerns the sliding block maxima estimator, but an analogous result can be derived for the disjoint block maxima estimator and under additional bias conditions for the circmax estimator. In view of the fact that the disjoint block maxima estimator exhibits a larger asymptotic estimation variance, the width of the disjoint blocks confidence interval is typically larger than the one of the circmax method, for every fixed confidence level (for an empirical illustration, see Figure 8 for the related problem of providing a confidence interval for the shape parameter $\gamma$).
6 Simulation study
The finite-sample properties of the circmax estimators and of the bootstrap approaches have been investigated in a large scale Monte-Carlo simulation study. Two asymptotic regimes were considered: first, the case where the block size $r$ is fixed and the number of disjoint blocks is increasing, and second, the case where $n$ is fixed and the block size is treated as a tuning parameter. All performance metrics in the subsequent paragraphs have been empirically approximated based on $N=5,000$ simulation repetitions.
6.1 Fixed block size
Throughout, we fix the block size $r$, and consider the estimation of target parameters associated with the law of $\bm{M}_{r}$, as well as bootstrap approximations for the respective estimation error. For illustrative purposes, we restrict attention to the one dimensional case and the most vanilla target parameter: the expected value $\mu_{r}=\operatorname{E}[M_{r}]$ of the block maximum distribution. The respective disjoint, sliding and circular block maxima estimators are $\hat{\mu}_{n}^{({\operatorname{mb}})}:=n^{-1}\sum_{i=1}^{n}M_{r,i}^{({\operatorname{mb}})}$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$, where ${\operatorname{cb}}(k)$ denotes the circmax estimator with parameter $k\in\{2,3\}$. The finite-sample performance is assessed for the following time series model.
Model 6.1 (ARMAX-GPD-Model).
$(X_{t})_{t\in\mathbb{Z}}$ is a stationary real-valued time series whose stationary distribution is the generalized Pareto distribution $\mathrm{GPD}(0,1,\gamma)$ with c.d.f.
$F_{\gamma}(x):=\begin{cases}\left(1-(1+\gamma x)^{-1/\gamma}\right)\mathds{1}(x\geq 0),\quad&\gamma>0,\\ \left(1-(1+\gamma x)^{-1/\gamma}\right)\mathds{1}(0\leq x\leq-1/\gamma),\quad&\gamma<0,\\ \left(1-\exp(-x)\right)\mathds{1}(x\geq 0),&\gamma=0.\end{cases}$
[…]
$Y_{t}=\max\big{(}\beta Y_{t-1},(1-\beta)W_{t}\big{)}\quad\forall t\in\mathbb{Z}$ (6.1)
[…]
Note in passing that $(X_{t})_{t}$ from Model 6.1 is exponentially beta mixing and satisfies Condition 2.1 with $Z\sim\mathrm{GEV}(\gamma)$, $a_{r}=\{r(1-\beta)\}^{\gamma}$ and $b_{r}=\{r(1-\beta)^{\gamma}-1\}/\gamma$. Moreover, the mean and variance of $M_{r}$ exist for $\gamma<1$ and $\gamma<1/2$, respectively, and in the latter case an application of Theorems 2.3 and 3.2 shows that, for $r=r_{n}\to\infty$ such that $r=o(n)$ and $\log(n)=o(r^{1/2})$,
$\frac{\sqrt{n/r}}{(r(1-\beta))^{\gamma}}\big{(}\hat{\mu}_{n}^{({\operatorname{mb}})}-\mu_{r}\big{)}\rightsquigarrow\mathcal{N}(0,\sigma^{2}_{{\operatorname{mb}}}),\quad{\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\},$
where the asymptotic variance is given by
$\sigma^{2}_{{\operatorname{db}}}=\begin{cases}\frac{g_{2}-g_{1}^{2}}{\gamma^{2}},&\gamma<1/2,\gamma\neq 0,\\ \frac{\pi^{2}}{6},&\gamma=0,\end{cases}\qquad\sigma^{2}_{{\operatorname{cb}}(k)}=\sigma^{2}_{{\operatorname{sb}}}=\begin{cases}4\Gamma(-2\gamma)I(\gamma),&\gamma<0,\\ 4(\log 4-1),&\gamma=0,\\ -\frac{g_{2}}{\gamma}I(\gamma)&0<\gamma<1/2.\end{cases}$ (6.2)
Here, $\Gamma$ denote the gamma function, $g_{j}=\Gamma(1-j\gamma)$ and $I(\gamma):=2\int_{0}^{1/2}\{\alpha_{2\gamma}(w)-1\}\{w^{-\gamma-1}(1-w)^{-\gamma-1}\}\,\mathrm{d}w$ with $\alpha_{c}(w)=w^{-1}\int_{0}^{w}(1-z)^{c}\,\mathrm{d}z$. Details are provided in Section A.
For the simulation experiments, the block size was fixed to $r=90$ as it roughly corresponds to the number of days within a season - a common block size in environmental applications. The effective sample size (i.e., the number of disjoint blocks of size $r$, abbreviated by $m=\lfloor n/r\rfloor$ hereafter) has been varied between $50$ and $100$ resulting in total sample sizes $n$ between $4,500$ and $9,000$
Performance of the estimators. We start by comparing the four estimators in terms of their variance, squared bias and mean squared error (MSE). For assessing the bias, the true first moment $\mu_{r}$ was approximated by a preliminary Monte Carlo simulation based on $10^{6}$ independent block maxima from Model 6.1 with block size $r=90.$ The results are summarized in Figure 2. As suggested by the theory, both the sliding blocks estimator and the two circular blocks estimators perform uniformly better than the respective disjoint blocks estimator. The improvement gets smaller as the observations are getting more heavy-tailed; a phenomenon that has already been observed in the literature; see e.g. Bücher and Zanger, (2023), Bücher and Staud, (2024). It is noteworthy that the additional bias effect introduced in the circular blocks estimators does not contribute in a relevant way; its contribution to the overall MSE was found to be at most $0.2\%$ and often much smaller. Furthermore, the serial dependence does not change the qualitative results significantly.
Figure 2: Mean estimation with fixed block size $r=90$. Top three rows: Mean squared error $\mathrm{MSE}(\hat{\mu}_{n}^{({\operatorname{mb}})})$. Bottom three rows: relative MSE with respect to the disjoint blocks method, i.e., $\mathrm{MSE}(\hat{\mu}_{n}^{({\operatorname{mb}})})/\mathrm{MSE}(\hat{\mu}_{n}^{({\operatorname{db}})})$.
Performance of the bootstrap. We consider each of the bootstrap estimators $\hat{\mu}_{n}^{({\operatorname{mb}}),\ast}$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$, with number of bootstrap replications set to $B=1,000$. Recall that our results from the previous sections imply that the conditional bootstrap estimation errors $\hat{\mu}_{n}^{({\operatorname{cb}}(2)),*}-\hat{\mu}_{n}^{({\operatorname{cb}}(2))}$ and $\hat{\mu}_{n}^{({\operatorname{cb}}(3)),*}-\hat{\mu}_{n}^{({\operatorname{cb}}(3))}$ are consistent for the estimation error $\hat{\mu}_{n}^{({\operatorname{sb}})}-\mu_{r}$, that $\hat{\mu}_{n}^{({\operatorname{db}}),*}-\hat{\mu}_{n}^{({\operatorname{db}})}$ is consistent for $\hat{\mu}_{n}^{({\operatorname{db}})}-\mu_{r}$ and that $\hat{\mu}_{n}^{({\operatorname{sb}}),*}-\hat{\mu}_{n}^{({\operatorname{sb}})}$ is inconsistent for $\hat{\mu}_{n}^{({\operatorname{sb}})}-\mu_{r}$. All four statements are illustrated in Figure 3 by means of QQ-plots and histograms for the parameter choice $\gamma=-0.2,\beta=0.5$ and $m=80$; other choices lead to similar results.
Figure 3: Estimation error $\hat{\mu}_{n}^{({\operatorname{mb}})}-\mu_{r}$ vs. bootstrap error $\hat{\mu}_{n}^{({\operatorname{mb}}^{\prime}),*}-\hat{\mu}_{n}^{({\operatorname{mb}}^{\prime})}$ for mean estimators with fixed block size $r=90$. Top row: qq-plots. Bottom row: histograms and kernel density estimates. Underlying data from Model 6.1 with $\gamma=-0.2,\beta=0.5$ and $m=80$.
In the next step, we evaluate the performance of the four bootstrap approaches in terms of their ability to provide accurate estimates of the estimation variance $\sigma_{\operatorname{mb}}^{2}(r)=\operatorname{Var}(\hat{\mu}_{n}^{({\operatorname{mb}})})$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\}$. For that purpose, for each fixed sample of size $n$, we estimate the respective variance by the empirical variance of the sample of bootstrap estimates; recall that each such sample is of size $B=1,000$, for every ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$. In Figure 4, we depict the average over the $N=5,000$ bootstrap estimates, along with the true parameters determined in a presimulation based on $10^{6}$ repetitions. We observe that the bootstrap estimates are close to the their target values for the disjoint method and for both circmax methods, while the naive sliding blocks bootstrap underestimates the true variance substantially, as mathematically demonstrated in Remark 4.3.
Figure 4: Bootstrap-based estimation of the estimation variance $\sigma_{\operatorname{mb}}^{2}(r)=\operatorname{Var}(\hat{\mu}_{n}^{({\operatorname{mb}})})$ with fixed sample size $n=1,000$. Left two columns: ${\operatorname{mb}}={\operatorname{sb}}$; the dashed line corresponds to $\sigma_{\operatorname{sb}}^{2}(r)$. Right two columns: ${\operatorname{mb}}={\operatorname{db}}$; the dashed line corresponds to $\sigma_{\operatorname{db}}^{2}(r)$.
Finally, we evaluate the performance of the bootstrap approaches in terms of their ability to provide accurate confidence intervals of pre-specified coverage; clearly, smaller intervals of similar coverage would be preferable. For that purpose, we restrict ourselves to the basic bootstrap confidence interval (Davison and Hinkley, , 1997) as illustrated in Corollary 5.6. The empirical coverage and the average widths of the respective intervals are depicted in Figure 5, where we omit the naive sliding method because of its inconsistency. We find that in most scenarios the desired coverage is almost reached by any method, in particular for lighter tail behavior. The disjoint blocks approach has the best coverage overall, albeit with only minimal advantages and sometimes on a par with or even slightly worse than the cb(2)-method. However, the price for this is universally wider confidence intervals.
Figure 5: Basic bootstrap confidence intervals for $\mu_{r}$ with fixed block size $r=90$. Top three rows: empirical coverage with intended coverage of $95\%$ (dashed line). Bottom three rows: relative average width with respect to the disjoint method, i.e., $\mathrm{width(CI}({\operatorname{db}}))/\mathrm{width(CI}({\operatorname{mb}}))$.
6.2 Fixed total sample size
In a second experiment, we consider the estimation of target parameters associated with the limiting law of $\bm{M}_{r}$ for $r\to\infty$, as well as bootstrap approximations for the respective estimation errors. For illustrative purposes, we restrict attention to the univariate, heavy tailed case and the estimation of the shape parameter $\alpha$ by pseudo-maximum likelihood estimation, as extensively discussed in Section 5.
Model 6.2 (ARMAX-Pareto-Model).
[…]
For the simulation experiment, we fix the total sample size to $n=1,000$, and treat the block size $r$ as a variable tuning parameter akin to the choice of $k$ in the peak-over-threshold method. More precisely, we consider choices of $r$ ranging from 8 to 40, resulting in effective sample sizes $m=n/r$ ranging from 125 to 25.
Performance of the estimators. In Figure 6 we depict the MSE of $\hat{\alpha}_{n}^{({\operatorname{mb}})}$ from Section 5 with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$ as a function of the effective sample size $m=n/r$. For $\beta=0.5$, we observe the typical bias-variance tradeoff resulting in a u-shaped MSE-curve: the larger the effective sample size, the smaller the variance and the larger the bias (for $\beta=0$, the MSE is dominated by the variance even for block sizes as small as 8). Apart from that, the findings are similar to the previous results for the case where $r$ was fixed: the sliding and circular block maxima estimator perform remarkably similar (in particular, the additional bias in the circular maxima is irrelevant), and they uniformly outperform the disjoint blocks estimator across all models under consideration.
Figure 6: Mean squared error for the estimation of $\alpha$ with fixed sample size $n=1,000$.
Performance of the bootstrap. Next, we consider each of the bootstrap estimators $\hat{\alpha}_{n}^{({\operatorname{mb}}),\ast}$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$, with number of bootstrap replications set to $B=1,000$. Similar as in Section 6.1, we start by evaluating the performance of the four bootstrap approaches in terms of their ability to provide accurate estimates of the estimation variance $\sigma_{\operatorname{mb}}^{2}(r)=\operatorname{Var}(\hat{\alpha}_{n}^{({\operatorname{mb}})})$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\}$. In Figure 7, we depict the average over the $N=5,000$ bootstrap estimates, along with the true parameters determined in a presimulation based on $10^{6}$ repetitions. The findings are akin to those in Section 6.1, the only additional remarkable observation being that the disjoint blocks method tends to overestimate the variance parameter in the iid case, in particular for large block sizes.
Figure 7: Bootstrap-based estimation of the estimation variance $\sigma_{\operatorname{mb}}^{2}(r)=\operatorname{Var}(\hat{\alpha}_{n}^{({\operatorname{mb}})})$ with fixed sample size $n=1,000$. Left two columns: ${\operatorname{mb}}={\operatorname{sb}}$; the dashed line corresponds to $\sigma_{\operatorname{sb}}^{2}(r)$. Right two columns: ${\operatorname{mb}}={\operatorname{db}}$; the dashed line corresponds to $\sigma_{\operatorname{db}}^{2}(r)$.
[…]
Figure 8: Basic bootstrap confidence intervals for $\alpha$ with fixed sample size $n=1,000$. Top three rows: empirical coverage with intended coverage of $95\%$ (dashed line). Bottom three rows: relative average width with respect to the disjoint method, i.e., $\mathrm{width(CI}({\operatorname{db}}))/\mathrm{width(CI}({\operatorname{mb}}))$.
7 Case study
[…]
Each moving window $w$ contains approximately $n=14,600$ daily observations, and the target parameter is estimated using the disjoint and sliding blocks estimator $\hat{\mu}_{n}^{({\operatorname{mb}})}=\hat{\mu}_{n}^{({\operatorname{mb}})}(w)$ with ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\}$. Basic bootstrap confidence intervals are constructed by either the classical approach based on resampling the disjoint block maxima (for estimator $\hat{\mu}_{n}^{({\operatorname{db}})}$), or by the circular block bootstrap approach with $k=2$ (for estimator $\hat{\mu}_{n}^{({\operatorname{sb}})}$). The resulting estimates and confidence bands are presented in Figure 9. The years on the x-axis correspond to the endpoints of the 40-year window. As expected from the theory (see in particular Figure 10), we find that the confidence intervals for the sliding blocks estimator are universally smaller than those for the disjoint blocks counterpart, with the average relative width being approximately 1.07.
Figure 9: Left plot: estimated expected annual maximum precipitation based on disjoint and sliding blocks estimators $\mu_{n}^{({\operatorname{db}})},\mu_{n}^{({\operatorname{sb}})}$ (solid lines), alongside with with $95\%$ bootstrap confidence intervals based on the db- and cb(2)-bootstrap, respectively (for a moving window of 40 years). Right plot: width of the respective confidence interval. All curves have been smoothed by a moving average filter of size 2.
8 Conclusion
Both the block-maxima method and the bootstrap are time-honored statistical methods that have seen wide use in applied statistics for extremes. Surprisingly, bootstrap consistency has never been proven, not even for the classical block-maxima method based on disjoint blocks. In this paper, respective consistency statements were established under high-level conditions on the data-generating process. A new approach, called the circular block-maxima method, has been proposed to allow for valid and computationally efficient bootstrap inference regarding the sliding block maxima method. The approach may be of independent interest for extreme-value analysis of non-stationary extremes, for instance in the presence of a temporal trend. Indeed, a reasonable model assumption inspired by many applications of the disjoint block maxima method would consist of the assumption that the circular block-maxima within a fixed $kr$-block of observations, say the $i$th one, follow an extreme-value distribution whose distribution/parameters are depending on $i$. Respective methods could be studied mathematically under a suitable triangular array structure, and we conjecture that advantages over the disjoint block maxima method will eventually show up in respective estimation variances. Other possible extensions would be the inclusion of inner-seasonal non-stationarities in the data-generating process or the generalization to empirical processes indexed by non-finite function classes.
9 Proofs
Proof of Theorem 3.2.
Throughout, we omit the upper index $({\operatorname{cb}}).$ We start by showing that
$\lim_{n\to\infty}\operatorname{Cov}(\bar{\mathbb{G}}_{n,r}\bm{h})=\Sigma_{\bm{h}}^{({\operatorname{sb}})}.$ (9.1)
By elementary arguments, it is sufficient to consider the case $q=1$. For $i\in\{1,\dots,m(k)\}$ write
$D_{n,i}=\frac{1}{kr}\sum_{s\in I_{kr,i}}\big{\{}h(\bm{Z}_{r,s})-\operatorname{E}[h(\bm{Z}_{r,s})]\big{\}},$ (9.2)
such that $\bar{\mathbb{G}}_{n,r}h_{j}=\sqrt{\frac{n}{r}}m(k)^{-1}\sum_{i=1}^{m(k)}D_{n,i}$. By stationarity and $n/r=km(k)$, we have
$\operatorname{Var}\big{(}\bar{\mathbb{G}}_{n,r}h\big{)}=k\operatorname{Var}(D_{n,1})+r_{n1}+r_{n2},$ (9.3)
where $r_{n1}=2k(1-\frac{1}{m(k)})\operatorname{Cov}(D_{n,1},D_{n,2})$ and $r_{n2}=2k\sum_{d=2}^{m(k)-1}(1-\frac{d}{m(k)})\operatorname{Cov}(D_{n,1},D_{n,1+d}).$ Thus, the proof of (9.1) is finished once we show that
$\lim_{n\to\infty}k\operatorname{Var}(D_{n,1})=\Sigma_{h}^{({\operatorname{sb}})},\quad\lim_{n\to\infty}r_{n1}=0,\quad\lim_{n\to\infty}r_{n2}=0.$ (9.4)
We start with the former, and for that purpose we define, for $\xi,\xi^{\prime}\in[0,k)$,
$f_{r}(\xi,\xi^{\prime})$ $:=\operatorname{Cov}\big{(}h(\bm{Z}_{r,1+\lfloor r\xi\rfloor}),h(\bm{Z}_{r,1+\lfloor r\xi^{\prime}\rfloor})\big{)},$
$g_{(k)}(\xi,\xi^{\prime})$ $:=\operatorname{Cov}\big{(}h(\bm{Z}^{(1)}_{|\xi-\xi^{\prime}|,(k)}),h(\bm{Z}^{(2)}_{|\xi-\xi^{\prime}|,(k)})\big{)},$
$g(\xi,\xi^{\prime})$ $:=\operatorname{Cov}\big{(}h(\bm{Z}^{(1)}_{|\xi-\xi^{\prime}|}),h(\bm{Z}^{(2)}_{|\xi-\xi^{\prime}|})\big{)},$
where $(\bm{Z}^{(1)}_{|\xi-\xi^{\prime}|,(k)},\bm{Z}^{(2)}_{|\xi-\xi^{\prime}|,(k)})$ has cdf $G_{|\xi-\xi^{\prime}|}^{(k)}$ from (11.1) and where $(\bm{Z}^{(1)}_{|\xi-\xi^{\prime}|},\bm{Z}^{(2)}_{|\xi-\xi^{\prime}|})$ has cdf $G_{|\xi-\xi^{\prime}|}$ from (2.3). Observe that, by Proposition 11.1, Condition 10.1 and Example 2.21 in van der Vaart, (1998), $\lim_{n\to\infty}f_{r}(\xi,\xi^{\prime})=g_{(k)}(\xi,\xi^{\prime})$. Hence, by Condition 10.1 and Dominated Convergence
$k\operatorname{Var}(D_{n,1})=\frac{1}{kr^{2}}\sum_{s=1}^{kr}\sum_{t=1}^{kr}\operatorname{Cov}\big{[}h(\bm{Z}_{r,s}),h(\bm{Z}_{r,t})\big{]}$ $=\frac{1}{k}\int_{0}^{k}\int_{0}^{k}f_{r}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi$
$=\frac{1}{k}\int_{0}^{k}\int_{0}^{k}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi+o(1).$ (9.5)
Now let $u(x):=g(x,0)$ and note that $g(\xi,\xi^{\prime})=u(|\xi-\xi^{\prime}|)$ and that $u(x)=0$ for $x>1$. By the definition of $G_{|\xi-\xi^{\prime}|}^{(k)}$ in (11.1) we have $G_{|\xi-\xi^{\prime}|}^{(k)}=G_{|\xi-\xi^{\prime}|}$ if $10\leq|\xi-\xi^{\prime}|\leq k-$ 1, which also implies that $g_{(k)}(\xi,\xi^{\prime})=g_{(k)}(|\xi-\xi^{\prime}|,0)=g(|\xi-\xi^{\prime}|,0)=u(|\xi-\xi^{\prime}|)$ for $\xi\leq k-1.$ Therefore,
$\int_{0}^{k}\int_{\xi}^{k}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi$ $=\int_{0}^{k}\int_{\xi}^{k\wedge(\xi+k-1)}u(\xi^{\prime}-\xi)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi+\int_{0}^{1}\int_{\xi+k-1}^{k}g_{(k)}(\xi^{\prime}-\xi,0)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi$
$=\int_{0}^{k}\int_{\xi}^{k\wedge(\xi+1)}u(\xi^{\prime}-\xi)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi+\int_{0}^{1}\int_{k-1+\xi}^{k}g_{(k)}(\xi^{\prime}-\xi,0)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi,$ (9.6)
where we used that $u(x)=0$ for $x>1$ at the second equality. After substituting $x=\xi^{\prime}-\xi$ in the inner integral and noting that $\int_{0}^{1}u(x)\,\mathrm{d}x=\Sigma_{h}^{({\operatorname{sb}})}/2,$ the first integral on the right-hand side of (9) can be written as
$\int_{0}^{k}\int_{\xi}^{k\wedge(\xi+1)}u(\xi^{\prime}-\xi)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=\int_{0}^{k}\int_{0}^{(k-\xi)\wedge 1}u(x)\,\mathrm{d}x\,\mathrm{d}\xi$ $=\int_{0}^{k-1}\int_{0}^{1}u(x)\,\mathrm{d}x\,\mathrm{d}\xi+\int_{k-1}^{k}\int_{0}^{k-\xi}u(x)\,\mathrm{d}x\,\mathrm{d}\xi$
$=(k-1)\frac{\Sigma_{h}^{({\operatorname{sb}})}}{2}+\int_{0}^{1}\int_{0}^{y}u(x)\,\mathrm{d}x\,\mathrm{d}y,$
where we used the substitution $y=k-\xi$ in the last step. Moreover, substituting $x=k-\xi^{\prime}+\xi$ in the inner integral, the second integral on the right-hand side of (9) can be written as
$\int_{0}^{1}\int_{k-1+\xi}^{k}g_{(k)}(\xi^{\prime}-\xi,0)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi$ $=\int_{0}^{1}\int_{\xi}^{1}g_{(k)}(k-x,0)\,\mathrm{d}x\,\mathrm{d}\xi=\int_{0}^{1}\int_{\xi}^{1}u(x)\,\mathrm{d}x\,\mathrm{d}\xi,$
where we used that $g_{(k)}(k-x,0)=u(x)$ for $x\in[0,1]$ by the definition of $G_{k-x}^{(k)}$. Overall, again using that $\int_{0}^{1}u(x)\,\mathrm{d}x=\Sigma_{h}^{({\operatorname{sb}})}/2$, the previous three displays imply that
$\int_{0}^{k}\int_{\xi}^{k}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=k\frac{\Sigma_{h}^{({\operatorname{sb}})}}{2}.$
This implies (9.4) in view of the symmetry $\int_{0}^{k}\int_{\xi}^{k}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=\int_{0}^{k}\int_{0}^{\xi}g_{(k)}(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi$ and (9).
Similar arguments as before, invoking asymptotic independence of $\bm{Z}_{r,s}$ and $\bm{Z}_{r,t}$ for $s\in I_{kr,1}$ and $t\in I_{kr,2}$ (see Proposition 11.1) and a similar identification of the covariances as integrals, imply that $r_{n1}=o(1)$. Finally, by Conditions 2.2(c) and 10.1 and Lemma 3.11 in Dehling and Philipp, (2002) we obtain
$\frac{|r_{n2}|}{k}$ $\lesssim m(k)\alpha(kr)^{\nu/(2+\nu)}\lesssim\Big{(}\Big{(}\frac{n}{r}\Big{)}^{1+2/\nu}\alpha(kr)\Big{)}^{v/(2+\nu)}=o(1),$
since $2/\omega<\nu$, where we have used that there is a lag of at least $kr$ between the observations making up $D_{n,1}$ and $D_{n,1+d}$. Overall, we have shown all three assertions in (9.4), and hence the proof of (9.1) is finished.
To verify the asymptotic normality, by the Cramér-Wold Theorem, it is again sufficient consider the case $q=1$. Recalling that $\bar{\mathbb{G}}_{n,r}h$ is centered, the case $\Sigma_{h}^{({\operatorname{sb}})}=0$ follows from (9.1); thus assume $\Sigma_{h}^{({\operatorname{sb}})}>0.$ Following Bücher and Segers, 2018a , we use a suitable blocking technique. Choose $m^{\ast}=m_{n}^{\ast}\in\mathbb{N}$ with $3\leq m^{\ast}\leq m(k)$ such that $m^{\ast}\to\infty$ and $m^{\ast}=o((n/r)^{\nu/(2(1+\nu))})$ with $\nu$ from Condition 10.1. Next, define $p:=p_{n}:=m(k)/m^{\ast}$ and assume for simplicity that $p\in\mathbb{N},n/r\in\mathbb{N}$. Furthermore, let $J_{j}^{+}:=\{(j-1)m^{\ast}+1,\ldots,jm^{\ast}-1\},J_{j}^{-}:=\{jm^{\ast}\}$ and note that $|J_{j}^{+}|=(m^{\ast}-1).$ Then,
$\bar{\mathbb{G}}_{n,r}h=\frac{1}{\sqrt{p}}\sum_{j=1}^{p}(S_{n,j}^{+}+S_{n,j}^{-}),\quad\text{ where }\quad S_{n,j}^{\pm}:=\sqrt{\frac{np}{r}}\frac{1}{m(k)}\sum_{i\in J_{j}^{\pm}}D_{n,i}$ (9.7)
with $D_{n,i}$ from (9.2). Noticing that the time series $(S_{n,j}^{-})_{j}$ is stationary, we have
[…]
By Conditions 2.2(a) and 10.1 we have $r_{n3}=O(1/m^{\ast})=o(1)$. Furthermore, by Condition 10.1 and Lemma 3.11 from Dehling and Philipp, (2002) with $s=r=1/(2+\nu)$ and $t=\nu/(2+\nu)$, we have $\sup_{d\geq 2}|\operatorname{Cov}(S_{n,1}^{-},S_{n,1+d}^{-})|\lesssim 10k(m^{\ast})^{-1}\alpha^{\nu/(2+\nu)}(kr)$, whence $r_{n4}\lesssim r^{-1}n(m^{\ast})^{-2}\alpha^{\nu/(2+\nu)}(kr)=o((m^{*})^{-2})=o(1)$ by Condition 2.2 and the fact that $2/\nu<\omega$. Put together, in view of $\operatorname{E}[S_{n,j}^{-}]=0$, these convergences imply $p^{-1/2}\sum_{j=1}^{p}S_{n,j}^{-}=o_{\mathbb{P}}(1).$
As there is a lag of at least $kr$ between the observations making up $S_{n,j_{1}}^{+}$ and $S_{n,j_{2}}^{+}$ for $j_{1}\neq j_{2}$ a standard argument involving characteristic functions, a complex-valued version of Lemma 3.9 in Dehling and Philipp, (2002), Condition 2.2 (b) and Levy’s Continuity Theorem allows for assuming that the $(S_{n,j}^{+})_{j}$ are independent. Hence, we may subsequently apply Lyapunov’s central limit theorem. To verify its conditions, note that, since $\operatorname{Var}(p^{-1/2}\sum_{j=1}^{p}S_{n,j}^{-})=o(1)$ and by using (9.1) and (9.7), $\operatorname{Var}(p^{-1/2}\sum_{j=1}^{p}S_{n,j}^{+})=\operatorname{Var}(\bar{\mathbb{G}}_{n,r}h)+o(1)=\Sigma_{h}^{({\operatorname{sb}})}+o(1).$ Hence,
$\frac{\sum_{j=1}^{p}\operatorname{E}[\big{|}S_{n,j}^{+}\big{|}^{2+\nu}]}{\{\operatorname{Var}(\sum_{j=1}^{p}S_{n,j}^{+})\}^{1+\nu/2}}\lesssim\frac{p(np/r)^{1+\nu/2}(m^{\ast}/m(k))^{2+\nu}}{(p\sigma^{2}/2)^{1+\nu/2}}\lesssim\frac{(m^{\ast})^{1+\nu}}{(n/r)^{\nu/2}}=o(1),$
by Condition 10.1 and the choice of $m^{\ast}.$ An application of Lyapunov’s central limit theorem implies the assertion.
To prove the additional weak convergences note that
$\tilde{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}h=\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}h+\sqrt{\frac{n}{r}}(P_{n,r}^{({\operatorname{cb}})}-P_{r})h,\quad\mathbb{G}_{n,r}^{({\operatorname{cb}})}h=\bar{\mathbb{G}}_{n,r}^{({\operatorname{cb}})}h+\sqrt{\frac{n}{r}}(P_{n,r}^{({\operatorname{cb}})}-P_{r}+P_{r}-P)h.$
Denote by $(\tilde{\bm{X}}_{t})_{t}$ an independent copy of $(\bm{X}_{t})_{t}$. By stationarity we have
$(P_{n,r}^{({\operatorname{cb}})}-P_{r})h=$ $\frac{1}{m(k)}\sum_{i=1}^{m(k)}\frac{1}{kr}\sum_{s=1}^{kr}\operatorname{E}[h(Z_{r,(i-1)kr+s})]-P_{r}h$
$=\frac{1}{kr}\sum_{s=(k-1)r+2}^{kr}\big{(}\operatorname{E}[h(Z_{r,s})]-P_{r}h\big{)}$
$=\frac{1}{kr}\sum_{t=1}^{r-1}\operatorname{E}\Big{[}h\Big{(}\frac{\max(X_{1},\dots,X_{t},X_{{(k-1)r+t+1}},\dots,X_{kr})-b_{r}}{a_{r}}\Big{)}-h\Big{(}\frac{\bm{M}_{r,1}-b_{r}}{a_{r}}\Big{)}\Big{]}.$
Condition 10.3 then gives $\sqrt{n/r}(P_{n,r}^{({\operatorname{cb}})}-P_{r})h\to k^{-1}(D_{h,k}+E_{h}).$ Furthermore, Condition 10.2 immediately implies $\sqrt{n/r}(P_{r}-P)h\to B_{h}.$ Combined with the first assertion and the penultimate display, we obtain the additional two claimed weak convergences. ∎
Proof of Theorem 4.1.
The second and third assertion follow immediately from the first one and the triangular inequality. For the proof of the first one, we omit the upper index ${\operatorname{cb}}$. By Theorem 2.3 (for ${\operatorname{mb}}={\operatorname{sb}}$) or Theorem 3.2 (for ${\operatorname{mb}}={\operatorname{cb}}$) and Lemma 2.3 from Bücher and Kojadinovic, (2019), it is sufficient to show that, unconditionally,
$(\hat{\mathbb{G}}_{n,r}^{[1]}\bm{h},\hat{\mathbb{G}}_{n,r}^{[2]}\bm{h})\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma^{({\operatorname{sb}})}_{h})\otimes\mathcal{N}_{q}(\bm{0},\Sigma^{({\operatorname{sb}})}_{h}),$
where, for $b\in\{1,2\},$ $\hat{\mathbb{G}}_{n,r}^{[b]}=\sqrt{\frac{n}{r}}\frac{1}{n}\sum_{i=1}^{m(k)}(W_{m(k),i}^{[b]}-1)\sum_{s\in I_{kr,i}}\delta_{\bm{Z}_{r,s}}$ with $\bm{W}^{[b]}_{m(k)}=(W_{m(k),1}^{[b]},\dots,W_{m(k),m(k)}^{[b]})$ two i.i.d. copies of $\bm{W}_{m(k)}$. By the Cramér-Wold device, it is sufficient to consider $q=1$, and for this in turn, yet again by the Cramér-Wold device, it is sufficient to show that
$T_{n}$ $:=\sum_{b=1}^{2}a_{b}\sqrt{\frac{n}{r}}\frac{1}{n}\sum_{i=1}^{m(k)}\big{(}W_{m(k),i}^{[b]}-1\big{)}\sum_{s\in I_{kr,i}}h(\bm{Z}_{r,s})\rightsquigarrow\mathcal{N}_{2}\big{(}0,(a_{1}^{2}+a_{2}^{2})\Sigma_{h}^{({\operatorname{sb}})}\big{)}$ (9.8)
for all $\bm{a}=(a_{1},a_{2})^{\top}\in\mathbb{R}^{2}$. Note that $\sum_{i=1}^{m(k)}\big{(}W_{m(k),i}^{[b]}-1\big{)}=0$ and that $\sum_{s\in I_{kr,i}}\operatorname{E}[h(\bm{Z}_{r,s})]$ does not depend on $i$, whence we may write
$T_{n}$ $=\sum_{b=1}^{2}a_{b}\sqrt{\frac{n}{r}}\frac{1}{n}\sum_{i=1}^{m(k)}\big{(}W_{m(k),i}^{[b]}-1\big{)}\Big{\{}\sum_{s\in I_{kr,i}}h(\bm{Z}_{r,s})-\operatorname{E}[h(\bm{Z}_{r,s})]\Big{\}}$
$=\sum_{b=1}^{2}a_{b}\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}\big{(}W_{m(k),i}^{[b]}-1\big{)}D_{n,i}$ (9.9)
with $D_{n,i}$ from (9.2). The subsequent proof strategy, known as Poissonization, consists of removing the dependence of the multinomial multipliers by introducing rowwise i.i.d. multiplier sequences that approximate the multinomial multipliers. More precisely, we employ the construction from Lemma 11.3: for each fixed $n\in\mathbb{N}$, let $(\bm{U}_{j,m(k)}^{[b]})_{j\in\mathbb{N},b\in\{1,2\}}$ be i.i.d. multinomial vectors with 1 trial and $m(k)$ classes, with class probabilities $1/m(k)$ for each class, and independent of $(\bm{X}_{t})_{t\in\mathbb{Z}}$. We may then assume that
${\bm{W}}_{m(k)}^{[b]}=\sum_{j=1}^{m(k)}\bm{U}_{j,m(k)}^{[b]}.$
Further, independent of $(\bm{U}_{j,m(k)}^{[b]})_{j\in\mathbb{N},b\in\{1,\dots,2\}}$ and of $(\bm{X}_{t})_{t\in\mathbb{Z}}$, let $(N_{m(k)}^{[b]})_{b=1,\ldots,M}$ be i.i.d. $\mathrm{Poisson}(m(k))$ distributed random variables. Define
$\tilde{\bm{W}}_{m(k)}^{[b]}=(\tilde{W}_{m(k),1}^{[b]},\dots,\tilde{W}_{m(k),m(k)}^{[b]})=\sum_{j=1}^{N_{m(k)}^{[b]}}\bm{U}_{j,m(k)}^{[b]}.$
By Lemma 11.3 the random vectors $\tilde{\bm{W}}_{m(k)}^{[1]},\tilde{\bm{W}}_{m(k)}^{[2]}$ are i.i.d. $\mathrm{Poisson}(1)^{\otimes m(k)}$ distributed. Then, in view of (9), we may write $T_{n}=\tilde{T}_{n}+a_{1}R_{n1}+a_{2}R_{n2}$, where
$\tilde{T}_{n}$ $=\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}\Big{\{}\sum_{b=1}^{2}a_{b}\big{(}\tilde{W}_{m(k),i}^{[b]}-1\big{)}\Big{\}}D_{n,i},$
$R_{nb}$ $=\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}\big{(}W_{m(k),i}^{[b]}-\tilde{W}_{m(k),i}^{[b]}\big{)}D_{n,i}.$
We may apply Lemma 11.2 to obtain that $\tilde{T}_{n}\rightsquigarrow\mathcal{N}(0,(a_{1}^{2}+a_{2}^{2})\Sigma_{h}^{({\operatorname{sb}})})$. For the proof of (9.8), it hence remains to verify that $R_{nb}=o_{\mathbb{P}}(1)$ for $b\in\{1,2\}$. We will suppress the upper index $[b]$ in the following. First write
$R_{nb}=\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}D_{n,i}\mathrm{sgn}(N_{m(k)}-m(k))\sum_{j=1}^{\infty}\bm{1}(i\in I_{m(k)}^{j}),$
where $I_{m(k)}^{j}=\{i\in\{1,\ldots,m(k)\}\colon|\tilde{W}_{m(k),i}-W_{m(k),i}|\geq j\}.$
We will start by showing that $\mathbb{P}(A_{n})=o(1),$ where $A_{n}:=\{|I_{m(k)}^{3}|>0\}=\{\exists i\in\{1,\dots,m(k)\}:|\tilde{W}_{m(k),i}-W_{m(k),i}|\geq 3\}$. Fix $\delta>0$. Invoking the central limit theorem, we may choose $C=C_{\delta}>0$ sufficiently large such that $\mathbb{P}(|N_{m(k)}-m(k)|>C\sqrt{m(k)})\leq\delta$ for all $n\in\mathbb{N}$. Next, note that, conditional on $N_{m(k)}=M$, the difference $|\tilde{W}_{m(k),i}-W_{m(k),i}|$ follows a $\mathrm{Bin}(|M-m(k)|,m(k)^{-1})$ distribution. Further, setting $c=\lceil C\sqrt{m(k)}\rceil$ and choosing $M$ such that $|M-m(k)|\leq c$, we have that $\mathrm{Bin}(|M-m(k)|,m(k)^{-1})([3,|M-m(k)|])\leq\mathrm{Bin}(c,m(k)^{-1})([3,c])$. As a consequence, conditioning on the event $N_{m(k)}=M$, we obtain
$\mathbb{P}(A_{n})$ $\leq\mathbb{P}(|I_{m(k)}^{3}|>0,|N_{m(k)}-m(k)|\leq c)+\delta$
$\leq\sum_{i=1}^{m(k)}\mathbb{P}(|\tilde{W}_{m(k),i}-W_{m(k),i}|\geq 3,|N_{m(k)}-m(k)|\leq c)+\delta$
$\leq m(k)\mathrm{Bin}(c,m(k)^{-1})([3,c])+\delta$
$\leq m(k)\{cm(k)^{-2}+\mathrm{Poisson}(cm(k)^{-1})([3,c])\}+\delta$
$=\frac{c}{m(k)}+m(k)e^{-c/m(k)}\sum_{j=3}^{c}\frac{(cm(k)^{-1})^{j}}{j!}+\delta$
$=O((m(k))^{-1/2})+\delta,$
where we used the approximation error of the Poisson Limit Theorem. Hence, since $\delta>0$ was arbitrary, $\mathbb{P}(A_{n})=o(1)$.
Next, note that, on $A_{n}^{c}$,
$|I_{m(k)}^{1}|+|I_{m(k)}^{2}|=\sum_{j=1}^{\infty}|I_{m(k)}^{j}|=\sum_{i=1}^{m(k)}\sum_{j=1}^{\infty}\bm{1}(i\in I_{m(k)}^{j})=\sum_{i=1}^{m(k)}|\tilde{W}_{m(k),i}-W_{m(k),i}|=|N_{m(k)}-m(k)|,$
using the fact that $\bm{U}_{j,m(k)}$ is multinomially distributed with 1 trial. As a consequence, letting $H_{n,i}=\bm{1}(N_{m(k)}\neq m(k))\{\bm{1}(i\in I_{m(k)}^{1})+\bm{1}(i\in I_{m(k)}^{2})\}/|N_{m(k)}-m(k)|$ and $\tilde{R}_{nb}=\big{|}\sum_{i=1}^{m(k)}H_{n,i}D_{n,i}\big{|},$ we obtain that
$|R_{nb}|$ $\leq\bm{1}_{A_{n}^{c}}\Big{|}\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}D_{n,i}\sum_{j=1}^{2}\bm{1}(i\in I_{m(k)}^{j})\Big{|}+o_{\mathbb{P}}(1)=\bm{1}_{A_{n}^{c}}\sqrt{k}\times\frac{|N_{m(k)}-m(k)|}{{\sqrt{m(k)}}}\times\tilde{R}_{nb}+o_{\mathbb{P}}(1).$
Therefore, since $|N_{m(k)}-m(k)|/\sqrt{m(k)}=O_{\mathbb{P}}(1)$ by the Central Limit Theorem, the proof of (9.8) is finished if we show that $\tilde{R}_{nb}=o_{\mathbb{P}}(1)$. For that purpose note that, for any fixed $\varepsilon>0$,
[…]
by the Central Limit Theorem, which in turn yields $\max_{i=1,\ldots,m(k)}H_{n,i}\leq 2|(N_{m(k)}-m(k))|^{-1}\bm{1}\{N_{m(k)}\neq m(k)\}=o_{\mathbb{P}}(1)$.
Now let $\sigma_{r}(i,j)=\operatorname{Cov}(D_{n,i},D_{n,j})$ and invoke the Conditional Tschebyscheff inequality to obtain that, for fixed $\varepsilon>0$,
[…]
$\lesssim\sum_{i,j=1}^{m(k)}H_{n,i}H_{n,j}|\sigma_{r}(i,j)|$
$\leq\Big{\{}\max_{j=1}^{m(k)}H_{n,j}\Big{\}}\times\Big{\{}\sum_{i=1}^{m(k)}H_{n,i}\sigma_{r}(1,1)+2\sum_{i=1}^{m(k)-1}H_{n,i}|\sigma_{r}(1,2)|+\sum_{|i-j|\geq 2}H_{n,i}|\sigma_{r}(i,j)|\Big{\}}$
$\leq\Big{\{}\max_{j=1}^{m(k)}H_{n,j}\Big{\}}\times\Big{\{}\sigma_{r}(1,1)+2|\sigma_{r}(1,2)|+2m(k)\max_{d=2}^{m(k)-1}|\sigma_{r}(1,1+d)|\Big{\}},$
where we used $\sum_{i=1}^{m(k)}H_{n,i}\leq 1$. Since $\max_{j=1}^{m(k)}H_{n,j}=o_{\mathbb{P}}(1)$ and by the (proof of the) three assertions in (9.4), we obtain that the expression on the right-hand side of the previous display is $o_{\mathbb{P}}(1)$. Hence, writing $\mathbb{P}(|\tilde{R}_{nb}|>\varepsilon)=\operatorname{E}\big{[}\mathbb{P}\big{(}|\tilde{R}_{nb}|>\varepsilon\mid(H_{n,i})_{i}\big{)}\big{]}$ and invoking the Dominated Convergence Theorem for convergence in probability, we obtain $\tilde{R}_{nb}=o_{\mathbb{P}}(1)$ and the proof is finished. ∎
Proof for Remark 4.3.
We sketch the proof of the conditional weak convergence in (4.2) provided that Conditions 2.1, 2.2 and  10.1(a) are met. Repeating the arguments from the proof of Theorem 4.1, the assertion follows from unconditional weak convergence of
$\tilde{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}h=\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}Y_{n,i}D_{n,i},$
where $(Y_{n,i})_{i}$ are iid with expectation 0 and variance 1, and where $D_{n,i}$ are defined as in (9.2), but with ${\operatorname{sb}}$ instead of ${\operatorname{cb}}$. We only provide the proof for the convergence of the variance of $\tilde{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}h$ to $\Sigma_{h}^{(k)}$; the remaining arguments are then the same as in the proof of Lemma 11.2.
Since $(Y_{n,i}D_{n,i})_{i}$ in uncorrelated, we obtain
$\operatorname{Var}(\tilde{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*)}h)=k\operatorname{Var}(D_{n,1})=\frac{1}{kr^{2}}\sum_{s=1}^{kr}\sum_{t=1}^{kr}\operatorname{Cov}(h(\bm{Z}_{r,s}^{({\operatorname{sb}})}),h(\bm{Z}_{r,t}^{({\operatorname{sb}})}))$
Using the notation after (9.4), we obtain
$\lim_{n\to\infty}\operatorname{Var}(\tilde{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*)}h)=\frac{1}{k}\int_{0}^{k}\int_{0}^{k}g(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi.$
We start by considering the case $k\geq 2$. Then, by the calculation in (9) but with $g_{(k)}$ replaced by $g$,
$\int_{0}^{k}\int_{\xi}^{k}g(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=I_{1}+I_{2}$
where
$I_{1}$ $=(k-1)\Sigma_{h}^{({\operatorname{sb}})}/2+\int_{0}^{1}\int_{0}^{y}u(x)\,\mathrm{d}x\,\mathrm{d}y,\qquad I_{2}=\int_{0}^{1}\int_{k-1+\xi}^{k}u(\xi^{\prime}-\xi)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=0.$
The first integral over $u$ can be rewritten as
$\int_{0}^{1}\int_{0}^{y}u(x)\,\mathrm{d}x\,\mathrm{d}y=\int_{0}^{1}\int_{x}^{1}u(x)\,\mathrm{d}y\,\mathrm{d}x=\int_{0}^{1}(1-x)u(x)\,\mathrm{d}x=\Sigma_{h}^{({\operatorname{sb}})}/2-\int_{0}^{1}xu(x)\,\mathrm{d}x,$ (9.10)
so that
$I_{1}=k\Sigma_{h}^{({\operatorname{sb}})}/2-\int_{0}^{1}xu(x)\,\mathrm{d}x.$
For symmetry reasons, we also have $\int_{0}^{k}\int_{0}^{\xi}g(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=\int_{0}^{k}\int_{\xi}^{k}g(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi$ so that
$\operatorname{Var}(\tilde{\mathbb{G}}_{n,r}^{({\operatorname{sb}}),*}h)=k\operatorname{Var}(D_{n,1})=\Sigma_{h}^{({\operatorname{sb}})}-\frac{2}{k}\int_{0}^{1}xu(x)\,\mathrm{d}x+o(1)$
as asserted.
It remains to consider the case $k=1$. A straightforward calculation then shows that
$\int_{0}^{1}\int_{\xi}^{1}g(\xi,\xi^{\prime})\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=\int_{0}^{1}\int_{\xi}^{1}u(\xi^{\prime}-\xi)\,\mathrm{d}\xi^{\prime}\,\mathrm{d}\xi=\int_{0}^{1}\int_{0}^{1-\xi}u(y)\,\mathrm{d}y\,\mathrm{d}\xi=\int_{0}^{1}\int_{0}^{x}u(y)\,\mathrm{d}y\,\mathrm{d}x,$
which implies the assertion by (9.10). ∎
Proof of Theorem 5.5.
The results regarding $(\hat{\alpha}_{n}^{({\operatorname{cb}})},\hat{\sigma}_{n}^{({\operatorname{cb}})})^{\top}$ and $(\hat{\alpha}_{n}^{({\operatorname{cb}}),*},\hat{\sigma}_{n}^{({\operatorname{cb}}),*})^{\top}$ follow from an application of Theorem 11.6 and Theorem 11.7, respectively.
For the former we need to show that, with $v_{n}=\sqrt{n/r},\omega_{n}=n$ and $X_{n,s}=M_{r,s}^{({\operatorname{cb}})}\vee c$, the “no-ties-condition” in Equation (11.4) is met, and that the three convergences in Condition 11.5 are satisfied. Equation (11.4) follows from $\mathbb{P}(X_{n,1}=\dots=X_{n,n})\leq\mathbb{P}(X_{n,1}=X_{n,r+1})=\mathbb{P}((M_{r,1}\vee c)/\sigma_{r}=(M_{r,r+1}\vee c)/\sigma_{r})$, which converges to zero by the Portmanteau theorem, observing that $(M_{r,1}\vee c)/\sigma_{r},(M_{r,r+1}\vee c)/\sigma_{r})$ weakly converges to the product of two independent Fréchet$(\alpha_{0},1)$ random variables by Lemma 5.1 in Bücher and Segers, 2018a . Condition 11.5(iii) is a consequence of Condition 5.4. Condition 11.5(ii) follows from a straightforward modification of Theorem 3.2 to the Fréchet domain of attraction. Finally, for $\alpha_{+}$ chosen sufficiently close to $\alpha_{0}$, Condition 11.5(ii) follows from such a modification as well, following the argumentation on page 1116 in Bücher and Segers, 2018b .
Next, consider the assertions regarding $(\hat{\alpha}_{n}^{({\operatorname{cb}}),*},\hat{\sigma}_{n}^{({\operatorname{cb}}),*})^{\top}$, which follow from Theorem 11.7 if we additionally show that the “no-ties-condition” in Equation (11.5) is met, that ^$\hat{\mathbb{P}}_{n,r}^{({\operatorname{cb}}),*}f=P_{\alpha_{0},1}f+o_{\mathbb{P}}(1)$ for all $f\in\mathcal{F}(\alpha_{-},\alpha_{+})$ and that (11.7) is met. The latter two assertions follow from similar arguments as given in the proof of Theorem 4.1 (adaptations to the Fréchet domain of attraction are needed), and it remains to show (11.5). For that purpose, let $I_{1},\dots,I_{m(k)}$ be iid uniformly distributed on $\{1,\dots,m(k)\}$ independent of $(X_{n,1},\ldots,X_{n,n})$, such that
$(X_{n,1}^{*},\dots,X_{n,n}^{*})\stackrel{{\scriptstyle d}}{{=}}(X_{n,kr(I_{1}-1)+1},\dots,X_{n,kr(I_{1}-1)+kr},\dots,X_{n,kr(I_{m(k)}-1)+1},\dots,X_{n,kr(I_{m(k)}-1)+kr}).$
Then
$\mathbb{P}(X_{n,1}^{\ast}=\ldots=X_{n,n}^{\ast})$ $=\sum_{i=1}^{m(k)}\mathbb{P}(I_{1}=i,X_{n,1}^{\ast}=\ldots=X_{n,n}^{\ast})$
$\leq\sum_{i=1}^{m(k)}\mathbb{P}(I_{1}=i,X_{n,kr(I_{1}-1)+1}=\ldots=X_{n,kr(I_{1}-1)+kr})$
$=\frac{1}{m(k)}\sum_{i=1}^{m(k)}\mathbb{P}(X_{n,kr(i-1)+1}=\ldots=X_{n,kr(I_{1}-1)+kr})=\mathbb{P}(X_{n,1}=\ldots=X_{n,kr}).$
This probability is bounded by $\mathbb{P}(X_{n,1}=X_{n,r+1})$, which converges to zero as shown above when proving Equation (11.4). ∎
Proof of Corollary 5.6.
Throughout, we omit the upper indexes $({\operatorname{sb}})$ and $({\operatorname{cb}})$ at $\hat{\sigma}_{n}$ and $\hat{\sigma}_{n}^{*}$, respectively. Define $S_{n}=\sqrt{\tfrac{n}{r}}(\hat{\sigma}_{n}/\sigma_{r_{n}}-1)$ and $S_{n}^{*}=\sqrt{\tfrac{n}{r}}(\hat{\sigma}_{n}^{*}-\hat{\sigma}_{n})/\sigma_{r_{n}}$, and note that, with $M_{2}(\alpha_{0})$ denoting the second row of $M(\alpha_{0})$,
$S_{n}=M_{2}(\alpha_{0})\mathbb{G}_{n,r}(f_{1},f_{2},f_{3})+o_{\mathbb{P}}(1),\qquad S_{n}^{*}=\frac{\hat{\sigma}_{n}}{\sigma_{r_{n}}}\sqrt{\frac{n}{r}}(\hat{\sigma}_{n}^{*}/\hat{\sigma}_{n}-1)=M_{2}(\alpha_{0})\hat{\mathbb{G}}_{n,r}^{*}(f_{1},f_{2},f_{3})+o_{\mathbb{P}}(1)$
by Theorem 5.5. Moreover, arguing as in the proof of that theorem, the weak limit of $S_{n}$ coincides with the conditional weak limit of $\hat{S}_{n}^{*}$ given the observations $\mathcal{X}_{n}$. Therefore, observing that $F_{S_{n}^{*}}(x)=\mathbb{P}(S_{n}^{*}\leq x\mid\mathcal{X}_{n})$ satisfies $F_{S_{n}^{*}}(x)=F_{\hat{\sigma}_{n}^{*}}(\hat{\sigma}_{n}+\sqrt{\tfrac{r}{n}}\cdot\sigma_{r_{n}}x)$ and hence
$\mathbb{P}(\sigma_{r_{n}}\in I_{n,\beta})$ $=\mathbb{P}\Big{[}\sqrt{\frac{n}{r}}\cdot\sigma_{r_{n}}^{-1}\{(F_{\hat{\sigma}_{n}^{*}})^{-1}(\tfrac{\beta}{2})-\hat{\sigma}_{n}\}\leq\sqrt{\frac{n}{r}}\Big{(}\frac{\hat{\sigma}_{n}}{\sigma_{r_{n}}}-1\Big{)}\leq\sqrt{\frac{n}{r}}\cdot\sigma_{r_{n}}^{-1}\{(F_{\hat{\sigma}_{n}^{*}})^{-1}(1-\tfrac{\beta}{2})-\hat{\sigma}_{n}\}\Big{]}$
$=\mathbb{P}\Big{[}(F_{S_{n}^{*}})^{-1}(\tfrac{\beta}{2})\leq S_{n}\leq(F_{S_{n}^{*}})^{-1}(1-\tfrac{\beta}{2})\Big{]},$
the assertion follows from Lemma 4.2 in Bücher and Kojadinovic, (2019). ∎
10 Additional conditions
In order to integrate $\bm{h}(\bm{Z}_{r,i})$ to the limit, we need mild asymptotic integrability conditions. For ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\}$ the condition is standard. As the circmax-sample permutates the underlying time series on a non-negligible part of the sample, we need a more involved assumption on asymptotic integrability. In most cases both conditions can be verified by similar arguments. Note that (b) implies (a) by letting $t=r$ in the supremum. Moreover, for $k=1$ (disjoint blocks case), (a) and (b) are equivalent.
Condition 10.1 (Asymptotic Integrability).
Fix $k\in\mathbb{N}.$ Let $\bm{h}=(h_{1},\ldots,h_{q})^{\top}\colon\mathbb{R}^{d}\to\mathbb{R}^{q}$ be an a.e. continuous function with respect to the Lebesgue-measure on $\mathbb{R}^{d}.$ There exists a $\nu>0$ such that:
  1. (a)
    $\limsup_{r\to\infty}\operatorname{E}[\|\bm{h}(\bm{Z}_{r,1})\|^{2+\nu}]<\infty.$
  2. (b)
    $\limsup_{r\to\infty}\sup_{t=1,\ldots,r}\operatorname{E}[\|\bm{h}\big{(}(\max\{\bm{X}_{1},\ldots,\bm{X}_{t},\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr}\}-\bm{b}_{r})/\bm{a}_{r}\big{)}\|^{2+\nu}]<\infty.$
Often we additionally impose that $\nu>2/\omega$ with $\omega$ from Condition 2.2.
The next two conditions specify the asymptotic bias resulting from the approximation of the distribution of the various block maxima of size $r_{n}$ by the extreme value distribution $G$.
Condition 10.2 (Disjoint and sliding blocks bias).
Let $r=r_{n}\to\infty$ with $r_{n}=o(n)$ and let $\bm{h}=(h_{1},\ldots,h_{q})^{\top}\colon\mathbb{R}^{d}\to\mathbb{R}^{q}$ be measurable such that $h(\bm{Z}_{r,1})$ and $h(\bm{Z})$ are integrable. The following limit exists:
$\bm{B}_{\bm{h}}=\lim_{n\to\infty}\sqrt{\frac{n}{r}}\big{\{}\operatorname{E}\big{[}\bm{h}(\bm{Z}_{r,1})]-\operatorname{E}\big{[}\bm{h}(\bm{Z})\big{]}\big{\}}.$ (10.1)
Circularization as applied within the cirmax-sample changes the order of observations used to calculate some of the cirular block maxima. The next conditions captures the resulting bias by decomposing it into two pieces, the first of which is due to a coupling with an independent copy $(\tilde{\bm{X}}_{t})_{t}$ of $(\bm{X}_{t})_{t}.$
Condition 10.3.
Let $r=r_{n}\to\infty$ with $r_{n}=o(n)$ and let $\bm{h}=(h_{1},\ldots,h_{q})^{\top}\colon\mathbb{R}^{d}\to\mathbb{R}^{q}$ be measurable and $k\in\mathbb{N}_{\geq 2}$ be fixed. The following expectations and limits exist:
$\begin{split}\bm{D}_{\bm{h},k}=&\lim_{n\to\infty}\sqrt{\frac{n}{r}}\cdot\frac{1}{r}\sum_{t=1}^{r-1}\operatorname{E}\Big{[}\bm{h}\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t},\bm{X}_{{(k-1)r+t+1}},\dots,\bm{X}_{kr})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\\ &\hskip 128.0374pt-\bm{h}\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t},\tilde{\bm{X}}_{1},\dots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\Big{]},\\ \bm{E}_{\bm{h}}=&\lim_{n\to\infty}\sqrt{\frac{n}{r}}\cdot\frac{1}{r}\sum_{t=1}^{r-1}\operatorname{E}\Big{[}\bm{h}\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t},\tilde{\bm{X}}_{1},\dots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}-\bm{h}\big{(}\bm{Z}_{r,1}\big{)}\Big{]}.\end{split}$ (10.2)
Note in passing that the asymptotic bias $\bm{E}_{\bm{h}}$ already appeared in Bücher and Zanger, (2023) in a slightly modified form in the setting of observing piecewise-stationary time series.
The following result implies that in many situations the asymptotic bias terms $D_{\bm{h},k}$ and $E_{\bm{h}}$ are negligible. The result should be understood as a proof of concept in the sense that one might formulate sharper results involving stronger integrability assumptions (e.g., boundedness of $h_{j}$) or a faster decay of mixing coefficients. Recall the beta-mixing coefficient for stationary time series, $\beta(\ell):=\beta(\sigma(\ldots,\bm{X}_{1},\bm{X}_{0}),\sigma(\bm{X}_{1+\ell},\bm{X}_{2+\ell},\ldots))$ with $\ell\in\mathbb{N}$, where $\beta(\mathcal{B},\mathcal{F})$ denotes the beta dependence coefficient between two sigma-fields $\mathcal{B}$ and $\mathcal{F}$ (Bradley, , 2005).
Lemma 10.4.
Fix $k\in\mathbb{N}_{\geq 2}$ and suppose that Condition 2.1 and 2.2 are met. Then $\bm{D}_{\bm{h},k}=\bm{E}_{\bm{h}}=0$, provided the time series $(\bm{X}_{t})_{t}$ is exponentially $\beta$-mixing (i.e., there exists $c>0,\lambda\in(0,1)$ such that $\beta(\ell)\leq c\lambda^{\ell}$ for all $\ell\in\mathbb{N}$), that $\log n=o(r^{1/2})$ and that Condition 10.1(b) and
$\limsup_{n\to\infty}\sup_{t=1,\ldots,r}\operatorname{E}\Big{[}\Big{\|}h\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t},\tilde{\bm{X}}_{1},\dots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\Big{\|}^{2+\nu}\Big{]}<\infty$ (10.3)
$\limsup_{n\to\infty}\sup_{t=\lfloor r^{1/2}\rfloor,\ldots,r}\operatorname{E}\Big{[}\Big{\|}h\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t-\lfloor r^{1/2}\rfloor},\tilde{\bm{X}}_{1},\dots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\Big{\|}^{2+\nu}\Big{]}<\infty$ (10.4)
$\limsup_{n\to\infty}\sup_{t=\lfloor r^{1/2}\rfloor,\ldots,r}\operatorname{E}\Big{[}\Big{\|}h\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t-\lfloor r^{1/2}\rfloor},\bm{X}_{t+1},\ldots,\bm{X}_{r})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\Big{\|}^{2+\nu}\Big{]}<\infty$ (10.5)
hold for some $\nu>2/\omega$ with $\omega$ from Condition 2.2, where $(\tilde{\bm{X}}_{t})_{t}$ is an independent copy of $(\bm{X}_{t})_{t}$.
Proof.
Without loss of generality we may assume that $q=1.$ First consider $D_{h,k}$ and define, for $t\in\{1,\ldots,r-1\}$,
$D_{n,h,k,t}=h\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t},\bm{X}_{{(k-1)r+t+1}},\dots,\bm{X}_{kr})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}-h\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t},\tilde{\bm{X}}_{1},\dots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}.$
Note that showing $\sup_{t=1,\ldots,r-1}(n/r)^{1/2}|\operatorname{E}[D_{n,h,k,t}]|=o(1)$ implies $D_{h,k}=0.$ By applying Berbee’s coupling Lemma (Berbee, , 1979) for each fixed $t\in\{1,\ldots,r-1\}$ to the vectors $(\bm{X}_{1},\ldots,\bm{X}_{t})$ and $(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr})$ we may assume that the random vector $(\tilde{\bm{X}}_{1},\ldots,\tilde{\bm{X}}_{r-t})$ satisfies $\mathbb{P}\big{(}(\tilde{\bm{X}}_{1},\ldots,\tilde{\bm{X}}_{r-t})\neq(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr})\big{)}\leq\beta(r)\leq c\lambda^{r};$ where the last inequality follows by assumption. Thus, by Hölder’s inequality and (10.3),
$\sup_{t=1,\ldots,r-1}\sqrt{\frac{n}{r}}\operatorname{E}[|D_{n,h,k,t}|]\lesssim\sqrt{\frac{n}{r}}\lambda^{r(1+\nu)/(2+\nu)}=\exp\Big{[}r\Big{(}\log\lambda\frac{1+\nu}{2+\nu}+\frac{\log(n/r)}{2r}\Big{)}\Big{]}.$
The expression on the right converges to zero since $\log n=o(r)$. As a consequence, $D_{h,k}=0$.
It remains to show that $E_{h}=0$. Writing
$E_{n,h,t}=h\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t},\tilde{\bm{X}}_{1},\dots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}-h(\bm{Z}_{r,1}).$
it is sufficient to show that $\sup_{t=1,\ldots,r-1}(n/r)^{-1/2}|\operatorname{E}[E_{n,h,t}]|=o(1)$. For that purpose, we split the supremum according to $t>r/2$ or $t\leq r/2$; both cases can then be treated similarly, and we only provide details on the former one. For simplicity, we assume that $r/2\in\mathbb{N}$.
Write $\ell=\lfloor r^{1/2}\rfloor$. The proof is finished once we show that
$\sup_{t=r/2,\dots,r}\Big{|}\sqrt{\frac{n}{r}}\operatorname{E}\Big{[}h\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t},\tilde{\bm{X}}_{1},\dots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}$
$\hskip 142.26378pt-h\Big{(}\frac{\max(\bm{X}_{1},\ldots,\bm{X}_{t-\ell},\tilde{\bm{X}}_{1},\ldots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\Big{]}\Big{|}=o(1)$ (10.6)
$\sup_{t=r/2,\dots,r}\Big{|}\sqrt{\frac{n}{r}}\operatorname{E}\Big{[}h\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t-\ell},\tilde{\bm{X}}_{1},\dots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}$
$\hskip 142.26378pt-h\Big{(}\frac{\max(\bm{X}_{1},\ldots,\bm{X}_{t-\ell},{\bm{X}}_{t+1},\ldots,{\bm{X}}_{r})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\Big{]}\Big{|}=o(1)$ (10.7)
$\sup_{t=r/2,\dots,r}\Big{|}\sqrt{\frac{n}{r}}\operatorname{E}\Big{[}h\Big{(}\frac{\max(\bm{X}_{1},\ldots,\bm{X}_{t-\ell},\bm{X}_{t+1},\ldots,\bm{X}_{r})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}-h(\bm{Z}_{r,1})\Big{]}\Big{|}=o(1).$ (10.8)
The proof of (10.7) is similar to the proof of $D_{h,k}=0$ given above (invoking (10.4) instead of (10.3)), and is therefore omitted (the final argument requires $\log n=o(r^{1/2})$, which is then exactly met by assumption). The proof of (10.7) is a simplified version of the proof of (10.8), so we only prove the latter for the sake of brevity. In view of Condition 10.1(b) and (10.5), an application of Hölder’s inequality implies
[…]
$\lesssim\sup_{t=r/2,\dots,r}\sqrt{\frac{n}{r}}\Big{\{}\mathbb{P}\big{(}\exists j\in\{1,\dots,d\}:\max(X_{t-\ell+1,j},\ldots,X_{t,j})>\max(X_{1,j},\ldots,X_{t-\ell,j},X_{t+1,j},\ldots,X_{r,j})\big{)}\Big{\}}^{\frac{1+\nu}{2+\nu}}.$
By the union-bound, it is sufficient to show that
$R_{n}$ $\equiv\sup_{t=r/2,\dots,r}\sqrt{\frac{n}{r}}\max_{j=1}^{d}\Big{\{}\mathbb{P}\big{(}\max(X_{t-\ell+1,j},\ldots,X_{t,j})>\max(X_{1,j},\ldots,X_{t-\ell,j},X_{t+1,j},\ldots,X_{r,j})\big{)}\Big{\}}^{\frac{1+\nu}{2+\nu}}$
$=o(1).$ (10.9)
Subsequently, we write $X_{t}$ instead of $X_{t-j}$, as all subsequent bounds are uniform in $j$. We then have
[…]
$\leq\sum_{i=t-\ell+1}^{t}\mathbb{P}\big{(}X_{i}>\max(X_{1},\ldots,X_{t-\ell},X_{t+1},\ldots,X_{r})\big{)}.$
For fixed $i$ in the previous sum, let $J\subset\{1,\dots,t-\ell\}$ denote the maximal set of indexes such that $|j_{1}-i|\geq\ell$ and $|j_{1}-j_{2}|\geq\ell$ for all distinct $j_{1},j_{2}\in J$; note that $|J|=O(r/\ell)$ since $t>r/2$. We then have
$\mathbb{P}\big{(}X_{i}>\max(X_{1},\ldots,X_{t-\ell},X_{t+1},\ldots,X_{r})\big{)}\leq\mathbb{P}\big{(}X_{i}>\max(X_{j}:j\in J)\big{)}.$
We may now successively apply Berbee’s coupling lemma (Berbee, , 1979) to construct a vector $(\check{X}_{j})_{j\in J}$ with iid coordinates that is independent of $X_{i}$ and satisfies $\check{X}_{j}=_{d}X_{j}$ for all $j\in J$ such that
$\mathbb{P}\big{(}X_{i}>\max(X_{s}:s\in J)\big{)}\leq\mathbb{P}\big{(}X_{i}>\max(\check{X}_{s}:s\in J)\big{)}+|J|\beta(\ell).$
More precisely, writing $J=\{j_{1},\dots,j_{|J|}\}$, the first application is to $X_{j_{1}}$ and $(X_{j_{2}},\dots,X_{j_{|J|}},X_{i})$. The second application is to $(\check{X}_{j_{1}},X_{j_{2}})$ and $(X_{j_{3}},\dots,X_{j_{|J|}},X_{i})$, where $\check{X}_{j_{1}}$ is the random variable constructed with the first application, and so on.
Since $\mathbb{P}\big{(}X_{i}>\max(\check{X}_{s}:s\in J)\big{)}=2^{-|J|}$ by Fubini’s theorem, we obtain from the last three displays that
$\sup_{t=r/2,\dots,r}\mathbb{P}\big{(}\max(X_{t-\ell+1},\ldots,X_{t})>\max(X_{1},\ldots,X_{t-\ell},X_{t+1},\ldots,X_{r})\big{)}\leq\ell 2^{-|J|}+\ell|J|\beta(r)\lesssim\ell 2^{-r/\ell}+r\lambda^{\ell}.$
In view of our choice of $\ell=\lfloor r^{1/2}\rfloor$, and letting $\zeta=\min(2,1/\lambda)>1$, the right hand side is bounded by $2r\zeta^{-\sqrt{r}}$, whence $R_{n}$ from (10) can be bounded by
$R_{n}\lesssim\sqrt{\frac{n}{r}}\big{(}r\zeta^{-\sqrt{r}}\big{)}^{(1+\nu)/(2+\nu)}=\exp\Big{\{}-\sqrt{r}\Big{(}\log(\zeta)\frac{1+\nu}{2+\nu}-\frac{\log r}{\sqrt{r}}\frac{1+\nu}{2+\nu}-\frac{\log(n/r)}{\sqrt{r}}\Big{)}\Big{\}}=o(1)$
by assumption on $r$. This proves (10), and the proof is finished. ∎
11 Auxiliary results
Proposition 11.1 (Joint weak convergence of circular block maxima).
Fix $k\in\mathbb{N}_{\geq 2}$ and suppose Conditions 2.1 and 2.2(a), (b) are met. Then, for any $\xi,\xi^{\prime}\in[0,k)$ and $\bm{x},\bm{y}\in\mathbb{R}^{d}$, we have
[…]
with $G_{\xi}$ from (2.3). Moreover, any circular block maxima taken from distinct $kr$-blocks are asymptotically independent, that is, for any $1\leq i\neq i^{\prime}\leq m(k),\xi,\xi^{\prime}\in[0,k)$ and $\bm{x},\bm{y}\in\mathbb{R}^{d}$,
$\lim_{n\to\infty}\mathbb{P}\Big{(}\bm{Z}_{r,(i-1)kr+1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\leq\bm{x},\bm{Z}_{r,(i^{\prime}-1)kr+1+\lfloor\xi^{\prime}r\rfloor}^{({\operatorname{cb}})}\leq\bm{y}\Big{)}=G_{1}(\bm{x},\bm{y})=G(\bm{x})G(\bm{y}).$ (11.2)
Proof.
We start by showing marginal convergence (i.e., $\bm{y}=\bm{\infty}$); the result then corresponds to Proposition 3.1. The case $\xi\leq k-1$ is trivial. For $\xi>k-1$ and fixed $\bm{x}\in\mathbb{R}^{d}$, we may proceed as in the proof of Lemma 2.4 in Bücher and Zanger, (2023) to obtain
$\mathbb{P}\Big{(}\bm{Z}_{r,1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\leq\bm{x}\Big{)}$ $=\mathbb{P}\Big{(}\frac{\max(\bm{X}_{1+\lfloor\xi r\rfloor},\ldots,\bm{X}_{kr})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{x},\frac{\max(\bm{X}_{1},\ldots,\bm{X}_{r+\lfloor\xi r\rfloor-kr})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{x}\Big{)}$
$=\mathbb{P}\Big{(}\frac{\max(\bm{X}_{1+\lfloor\xi r\rfloor},\ldots,\bm{X}_{kr})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{x}\Big{)}\mathbb{P}\Big{(}\frac{\max(\bm{X}_{1},\ldots,\bm{X}_{r+\lfloor\xi r\rfloor-kr})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{x}\Big{)}+R_{n},$
where $R_{n}=O(\alpha(r_{n}))=o(1)$. Next, by stationarity, the product on the right-hand side of the previous display can be written as
$\mathbb{P}\Big{(}\bm{Z}_{kr-\lfloor\xi r\rfloor}\leq\frac{\bm{a}_{r}\bm{x}+\bm{b}_{r}-\bm{b}_{kr-\lfloor\xi r\rfloor}}{\bm{a}_{kr-\lfloor\xi r\rfloor}}\Big{)}\mathbb{P}\Big{(}\bm{Z}_{r+\lfloor\xi r\rfloor-kr}\leq\frac{\bm{a}_{r}\bm{x}+\bm{b}_{r}-\bm{b}_{r+\lfloor\xi k\rfloor-kr}}{\bm{a}_{r+\lfloor\xi r\rfloor-kr}}\Big{)}$
The convergence in Equation (2.1) being locally uniform we obtain, for $j\in\{1,\dots,d\}$,
$\lim_{n\to\infty}\frac{a_{r}^{(j)}x^{(j)}+b_{r}^{(j)}-b^{(j)}_{kr-\lfloor\xi r\rfloor}}{a^{(j)}_{kr-\lfloor\xi r\rfloor}}=(k-\xi)^{-\gamma^{(j)}}x^{(j)}+\frac{(k-\xi)^{-\gamma^{(j)}}-1}{\gamma^{(j)}},$
$\lim_{n\to\infty}\frac{a_{r}^{(j)}x^{(j)}+b_{r}^{(j)}-b^{(j)}_{r+\lfloor\xi r\rfloor-kr}}{a^{(j)}_{r+\lfloor\xi r\rfloor-kr}}=(1+\xi-k)^{-\gamma^{(j)}}x^{(j)}+\frac{(1+\xi-k)^{-\gamma^{(j)}}-1}{\gamma^{(j)}}.$
Arguing as in the proof of Lemma 8.9 from Bücher and Staud, (2024), Condition 2.1 and the previous three displays yield, for $n\to\infty,$
$\mathbb{P}(\bm{Z}_{r,1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\leq\bm{x})$ $=G^{k-\xi}(\bm{x})G^{1+\xi-k}(\bm{x})+o(1)=G(\bm{x})+o(1).$
We proceed by proving (11.1), and start by considering the case $|\xi-\xi^{\prime}|>k-1$. Without loss of generality let $\xi<\xi^{\prime}$ such that $1+\lfloor\xi r\rfloor\leq r$ and $1+\lfloor\xi^{\prime}r\rfloor>(k-1)r$. Then, by the same arguments as for the marginal convergence,
[…]
$=\lim_{n\to\infty}\mathbb{P}\Big{(}\frac{\max(\bm{X}_{r,1},\ldots,\bm{X}_{r,\lfloor\xi r\rfloor})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{y},\frac{\max(\bm{X}_{r,1+\lfloor\xi r\rfloor},\ldots,\bm{X}_{r,\lfloor\xi^{\prime}r\rfloor-(k-1)r})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{x}\wedge\bm{y}$
$\hskip 85.35826pt\frac{\max(\bm{X}_{r,\lfloor\xi^{\prime}r\rfloor-(k-1)r+1},\ldots,\bm{X}_{r,r+\lfloor\xi r\rfloor})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{x},\frac{\max(\bm{X}_{r,1+\lfloor\xi^{\prime}r\rfloor},\ldots,\bm{X}_{kr})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{y}\Big{)}$
$=G^{\xi}(\bm{y})\cdot G^{\xi^{\prime}-(k-1)-\xi}(\bm{x}\wedge\bm{y})\cdot G^{1+\xi-\xi^{\prime}+(k-1)}(\bm{x})\cdot G^{k-\xi^{\prime}}(\bm{y})$
$=G^{k-(\xi^{\prime}-\xi)}(\bm{y})\cdot G^{1-(k-(\xi^{\prime}-\xi))}(\bm{x}\wedge\bm{y})\cdot G^{k-(\xi^{\prime}-\xi)}(\bm{x})$
$=G_{k-(\xi^{\prime}-\xi)}(\bm{x},\bm{y}),$
where we used (2.3) at the last equality.
Next, consider the case $|\xi-\xi^{\prime}|\leq k-1$, and again assume $\xi<\xi^{\prime}$. There are two subcases to handle; first let $\xi^{\prime}\in(k-1,k].$ In that case we have $\xi\geq 1,$ hence there is no overlapping from the left (induced by the $\operatorname{circ{\text{-}}max}$-operation). We may then proceed in a similar (but simpler) way as for $|\xi-\xi^{\prime}|>k-1$ to obtain the stated limit. Alternatively, we have $\xi^{\prime}\leq k-1.$ In that case, we are in the sliding block maxima case, where the result is known, see e.g. Lemma 8.6 in Bücher and Staud, (2024).
Finally, the assertion in (11.2) follows from similar arguments as in Lemmas A.7, A.8 from Bücher and Segers, 2018b , which yield
$\mathbb{P}\big{(}\bm{Z}_{(i-1)kr+1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\leq\bm{x},\bm{Z}_{(i^{\prime}-1)kr+1+\lfloor\xi^{\prime}r\rfloor}^{({\operatorname{cb}})}\leq\bm{y}\big{)}$ $=\mathbb{P}\big{(}\bm{Z}_{1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\leq\bm{x}\big{)}\cdot\mathbb{P}\big{(}\bm{Z}_{1+\lfloor\xi^{\prime}r\rfloor}^{({\operatorname{cb}})}\leq\bm{y}\big{)}+o(1)$
by stationarity. The product on the right-hand side converges to $G(\bm{x})G(\bm{y})$ by the marginal convergence. ∎
Lemma 11.2 (Multiplier central limit theorem).
Suppose Conditions 2.1 and 2.2 are met. Fix $k\in\mathbb{N}$ and let $\bm{h}=(h_{1},\dots,h_{q})^{\top}$ satisfy Condition 10.1 with $\nu>2/\omega$ and $\omega$ from Condition 2.2. Let $Y_{n,1},\ldots,Y_{n,m(k)}$ be a triangular array of rowwise independent and identically distributed random variables that is independent of $(\bm{X}_{1},\bm{X}_{2},\dots)$ such that $\limsup_{n\to\infty}\operatorname{E}[|Y_{n,1}|^{2+\nu}]<\infty$ with $\nu$ from Condition 10.1 and such that the limit $\mu_{2}:=\lim_{n\to\infty}\operatorname{E}[Y_{n,1}^{2}]>0$ exists. Then, for $n\to\infty$ and with $\Sigma_{\bm{h}}^{({\operatorname{mb}})}$ from (2.4), we have
$\bar{\mathbb{G}}_{n,r}^{\circ}\bm{h}:=\sqrt{\frac{n}{r}}\frac{1}{n}\sum_{i=1}^{m(k)}Y_{n,i}\sum_{s\in I_{kr,i}}\big{\{}\bm{h}(\bm{Z}_{r,s}^{({\operatorname{cb}})})-\operatorname{E}[\bm{h}(\bm{Z}_{r,s}^{({\operatorname{cb}})})]\big{\}}\rightsquigarrow\begin{cases}\mathcal{N}_{q}(0,\mu_{2}\Sigma_{\bm{h}}^{({\operatorname{db}})})&\text{ if }k=1,\\ \mathcal{N}_{q}(0,\mu_{2}\Sigma_{\bm{h}}^{({\operatorname{sb}})})&\text{ if }k\in\{2,3,\dots\}.\end{cases}$
Proof.
We only consider the case $k\geq 2$; the case $k=1$ is a simple modification. By the Cramér-Wold Theorem, it is sufficient to consider the case $q=1$. Since $m(k)=n/(kr)$, we may write $\bar{\mathbb{G}}_{n,r}^{\circ}h=\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}Y_{n,i}D_{n,i}$, with $D_{n,i}=\frac{1}{kr}\sum_{s\in I_{kr,i}}\{h(\bm{Z}_{r,s}^{({\operatorname{cb}})})-\operatorname{E}[h(\bm{Z}_{r,s}^{({\operatorname{cb}})})]\}$ from (9.2). Then, since $(Y_{n,i}D_{n,i})_{i}$ is uncorrelated,
$\operatorname{Var}(\bar{\mathbb{G}}_{n,r}^{\circ}h)$ $=k\operatorname{E}[Y_{n,1}^{2}]\operatorname{Var}(D_{n,1})=\mu_{2}\Sigma_{h}^{({\operatorname{sb}})}+o(1)$
by (9.4). This implies the assertion for $\Sigma_{h}^{({\operatorname{sb}})}=0$.
If $\Sigma_{h}^{({\operatorname{sb}})}>0$, we may follow a similar line of reasoning as in the proof of Theorem 3.2: decompose
$\bar{\mathbb{G}}_{n,r}^{\circ}h=\frac{1}{\sqrt{p}}\sum_{j=1}^{p}(T_{n,j}^{+}+T_{n,j}^{-}),\quad\text{ where }\quad T_{n,j}^{\pm}:=\sqrt{\frac{np}{r}}\frac{1}{m(k)}\sum_{i\in J_{j}^{\pm}}Y_{n,i}D_{n,i},$
where $p=m(k)/m^{\ast}$ with $m^{\ast}=o(m^{\nu/(2(1+\nu))})$ and $J_{j}^{+}:=\{(j-1)m^{\ast}+1,\ldots,jm^{\ast}-1\},J_{j}^{-}:=\{jm^{\ast}\}$. Following the arguments from the proof of Theorem 3.2, we have $p^{-1/2}\sum_{j=1}^{p}T_{n,j}^{-}=o_{\mathbb{P}}(1)$. Moreover, we may assume independence of $(S_{n,j}^{+})_{j}$, which enables an application of Ljapunov’s central limit theorem to $p^{-1/2}\sum_{j=1}^{p}T_{n,j}^{+}$ to conclude. ∎
The next result is possibly well-known; we skip its elementary proof (see also Section 2 in Klaassen and Wellner, , 1992).
Lemma 11.3 (Poissonization).
For $q\in\mathbb{N}_{\geq 2}$ and $(p_{1},\dots,p_{q})\in[0,1]^{q}$ with $p_{1}+\dots+p_{q}=1$, let $(\bm{U}_{j,q})_{j\in\mathbb{N}}$ be i.i.d. multinomial vectors with $1$ trial and $q$ classes with class probabilities $p_{1},\dots,p_{q}$. For $m\in\mathbb{N}$ let $\bm{W}_{m,q}=(W_{m,q,1},\dots,W_{m,q,q})=\sum_{j=1}^{m}\bm{U}_{j,q}$ (which is multinomial with $m$ trials and $q$ classes). Further, for $\lambda>0$, let $N$ be a $\mathrm{Poi}(\lambda)$-distributed random variable that is independent of $(\bm{U}_{j,q})_{j\in\mathbb{N}}$. Then, the random variables $W_{N,q,1},\dots,W_{N,q,q}$ are independent and $W_{N,q,j}$ follows a $\mathrm{Poi}(\lambda p_{j})$-distribution for $j=1,\dots,q$.
Lemma 11.4.
Let $d=1$ and suppose that $h$ is a real-valued function such that $\operatorname{Var}(h(\bm{Z}))\in(0,\infty)$ exists. Then, for all $\xi\in[0,1]$, $\operatorname{Cov}(h(\bm{Z}_{1,\xi}),h(\bm{Z}_{2,\xi}))\geq 0$ with strict inequality in a neighbourhood of 00.
Proof.
By our assumption $d=1$, the cdf $G$ from (2.2) corresponds to the $\mathrm{GEV}(\gamma)$- distribution, which has a Lebesgue-density that we denote by $g$. For $\xi\in(0,1)$, we may employ the following stochastic construction for $(\bm{Z}_{1,\xi},\bm{Z}_{2,\xi})^{\top}$: let $U,W$ be i.i.d. with cdf $G^{\xi}$ and independent of $V,$ the latter having the cdf $G^{1-\xi}.$ We then have $\mathcal{L}(\bm{Z}_{1,\xi},\bm{Z}_{2,\xi})=\mathcal{L}(U\vee V,W\vee V).$ Without loss of generality assume $\operatorname{E}[h(\bm{Z}_{1,\xi})]=0$ and note that
$\operatorname{Cov}(h(\bm{Z}_{1,\xi}),h(\bm{Z}_{2,\xi}))$ $=\operatorname{E}[h(U\vee V)h(W\vee V)]$
$=2\operatorname{E}[h(U)h(V)G^{\xi}(V)\bm{1}(V<U)]$
$\hskip 28.45274pt+\operatorname{E}[h(U)h(W)G^{1-\xi}(U\wedge W)]+\operatorname{E}[h^{2}(V)G^{2\xi}(V)]\equiv 2I_{1}+I_{2}+I_{3}.$ (11.3)
Let $H\colon\bar{\mathbb{R}}\to\mathbb{R},H(x)=\int_{-\infty}^{x}h(z)g(z)\,\mathrm{d}z$ and note that $H(\infty)=0=H(-\infty)$, where we made use of $\operatorname{E}[h(\bm{Z}_{1,\xi})]=\int_{\mathbb{R}}h(z)g(z)\,\mathrm{d}z$. Using integration by parts we obtain
$I_{1}=(1-\xi)\xi\int_{-\infty}^{\infty}G^{\xi-1}(x)H^{\prime}(x)H(x)\,\mathrm{d}x=\frac{\xi}{2}(1-\xi)^{2}\int_{-\infty}^{\infty}G^{\xi-2}(x)g(x)H^{2}(x)\,\mathrm{d}x\geq 0.$
Similarly, we have $I_{2}=\xi^{2}(1-\xi)\int_{-\infty}^{\infty}G^{\xi-2}(x)g(x)H^{2}(x)\,\mathrm{d}x\geq 0.$ In view of (11) and $I_{3}\geq 0$ this proves the first assertion.
For the second assertion note that $I_{3}=(1-\xi)\operatorname{E}[h^{2}(\bm{Z})G^{\xi}(\bm{Z})]\leq\operatorname{Var}(h(\bm{Z}))<\infty$ and $I_{j}\geq 0$ for $j=1,2,3$. By dominated convergence and (11) we then have
$\liminf_{\xi\downarrow 0}\operatorname{Cov}(h(\bm{W}_{1,\xi}),h(\bm{W}_{2,\xi}))\geq\liminf_{\xi\downarrow 0}I_{3}=\operatorname{Var}(h(\bm{Z}))>0,$
which lets us conclude. ∎
11.1 A general result on bootstrapping the Fréchet independence MLE
Throughout this section, let $\mathcal{X}_{n}=(X_{n,1},\dots,X_{n,\omega_{n}})$ denote a sequence of random vectors in $(0,\infty)^{\omega_{n}}$ with continuous cumulative distribution functions $F_{n,1},\dots,F_{n,\omega_{n}}$, where $\omega_{n}\to\infty$ is a sequence of integers. Throughout, we assume that asymptotically not all observations are tied, that is,
$\lim_{n\to\infty}\mathbb{P}(X_{n,1}=\dots=X_{n,\omega_{n}})=0.$ (11.4)
For some $\alpha_{0}>0$ and $\sigma_{n}\to\infty$, proximity of the $F_{n,1},\dots,F_{n,\omega_{n}}$ to the Fréchet-distribution $P_{\alpha_{0},\sigma_{n}}$ will be controlled by convergence conditions on empirical means of $X_{n,i}/\sigma_{n}$: for $0<\alpha_{-}<\alpha_{0}<\alpha_{+}<\infty$, let $\mathcal{F}(\alpha_{-},\alpha_{+})$ denote the set of functions containing $x\mapsto\log x,x\mapsto x^{-\alpha},x\mapsto x^{-\alpha}\log x$ and $x\mapsto x^{-\alpha}\log^{2}x$ for all $\alpha\in(\alpha_{-},\alpha_{+})$. Moreover, recall $f_{1}(x)=x^{-\alpha_{0}}$, $f_{2}(x)=x^{-\alpha_{0}}\log x$ and $f_{3}(x)=\log x$ from (5.4), considered as functions on $(0,\infty)$, and let =$\mathcal{H}=\{f_{1},f_{2},f_{3}\}$. For a real-valued function $f$ defined on $(0,\infty)$ such that the following integrals/expectations exist, let
$\mathbb{P}_{n}f=\frac{1}{\omega_{n}}\sum_{i=1}^{\omega_{n}}f(X_{n,i}/\sigma_{n}),\quad P_{n}f=\frac{1}{\omega_{n}}\sum_{i=1}^{\omega_{n}}\operatorname{E}[f(X_{n,i}/\sigma_{n})],\quad P_{\alpha_{0},1}f:=\int f(x)\,\mathrm{d}P_{\alpha_{0},1}(x).$
Condition 11.5.
  1. (i)
    There exists $0<\alpha_{-}<\alpha_{0}<\alpha_{+}<\infty$ such that $\mathbb{P}_{n}f\rightsquigarrow P_{\alpha_{0},1}f$ for all $f\in\mathcal{F}(\alpha_{-},\alpha_{+})$.
  2. (ii)
    There exists $0<v_{n}\to\infty$ such that $\bar{\mathbb{G}}_{n}(f_{1},f_{2},f_{3})^{\top}\rightsquigarrow N_{3}(\bm{0},\Sigma_{G})$, where $\bar{\mathbb{G}}_{n}f=v_{n}(\mathbb{P}_{n}f-P_{n}f)$ and where $\Sigma_{G}\in\mathbb{R}^{3\times 3}$ is positive semidefinite.
  3. (iii)
    With $v_{n}$ from (ii), the limit $\bm{B}_{G}:=\lim_{n\to\infty}B_{n}(f_{1},f_{2},f_{3})^{\top}$ exists, where $B_{n}f=v_{n}(P_{n}f-P_{\alpha_{0},1}f)$.
Recall the log-likelihood function $\ell_{\bm{\theta}}$ of the Fréchet distribution from (5.2).
Theorem 11.6 (Bücher and Segers, 2018b, ).
(i) Suppose that (11.4) holds and that Condition 11.5(i) is met. Then, on the complement of the event $\{X_{n,1}=\dots=X_{n,\omega_{n}}\}$, the independence Fréchet log-likelihood ${\bm{\theta}}\mapsto\sum_{i=1}^{\omega_{n}}\ell_{\bm{\theta}}(X_{n,i})$ has a unit maximizer $\hat{\bm{\theta}}_{n}=(\hat{\alpha}_{n},\hat{\sigma}_{n})$, and that maximizer is consistent in the sense that $(\hat{\alpha}_{n},\hat{\sigma}_{n}/\sigma_{n})=(\alpha_{0},1)+o_{\mathbb{P}}(1)$ as $n\to\infty$.
(ii) If, additionally, Condition 11.5(ii)-(iii) is met, then
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}-\alpha_{0}\\ \hat{\sigma}_{n}/\sigma_{n}-1\end{pmatrix}=M(\alpha_{0})(\bar{\mathbb{G}}_{n}+B_{n})(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)\rightsquigarrow\mathcal{N}_{2}(M(\alpha_{0})\bm{B}_{G},M(\alpha_{0})\Sigma_{G}M(\alpha_{0})^{\top})$
as $n\to\infty$, where
$M(\alpha_{0})=\frac{6}{\pi^{2}}\begin{pmatrix}\alpha_{0}^{2}&\alpha_{0}(1-\gamma)&-\alpha_{0}^{2}\\ \gamma-1&-(\Gamma^{\prime\prime}(2)+1)/\alpha_{0}&1-\gamma\end{pmatrix}$
with $\Gamma(z)=\int_{0}^{\infty}t^{z-1}{e}^{-t}\,\mathrm{d}t$ the Euler Gamma function and $\gamma=0.5772\ldots$ the Euler-Mascheroni constant.
Now, conditional on $\mathcal{X}_{n}=(X_{n,1},\dots,X_{n,\omega_{n}})$, let $\mathcal{X}_{n}^{*}=(X_{n,1}^{*},\dots,X_{n,\omega_{n}}^{*})$ denote a bootstrap sample of $\mathcal{X}_{n}$; formally, $\mathcal{X}_{n}^{*}$ is assumed to be a measurable function of both $\mathcal{X}_{n}$ and of some additional independent random element $\mathcal{W}_{n}$ taking values in some measurable space. For the subsequent consistency statements, the bootstrap sample is assumed to satisfy the not-all-tied condition
$\lim_{n\to\infty}\mathbb{P}(X_{n,1}^{*}=\dots=X_{n,\omega_{n}}^{*})=0$ (11.5)
or, equivalently, $\mathbb{P}\big{(}X_{n,1}^{*}=\dots=X_{n,\omega_{n}}^{*}\mid\mathcal{X}_{n}\big{)}=o_{\mathbb{P}}(1)$ as $n\to\infty$. The bootstrap scheme is assumed to be regular in the sense that the conditional distribution of certain rescaled arithmetic means is close to the distribution of respective arithmetic means of the original sample. For a real-valued function $f$ on $(0,\infty)$, define ^$\hat{\mathbb{P}}_{n}^{*}f=\frac{1}{\omega_{n}}\sum_{i=1}^{\omega_{n}}f(X_{n,i}^{*}/\sigma_{n})$ and $\hat{\mathbb{G}}_{n}^{*}f=v_{n}(\hat{\mathbb{P}}_{n}^{*}f-\mathbb{P}_{n}f)$.
Theorem 11.7.
Suppose that (11.4), (11.5) and Condition 11.5 is met, that ^$\hat{\mathbb{P}}_{n}^{*}f=P_{\alpha_{0},1}f+o_{\mathbb{P}}(1)$ for all $f\in\mathcal{F}(\alpha_{-},\alpha_{+})$ (or, equivalently, (|^$\mathbb{P}(|\hat{\mathbb{P}}_{n}^{*}f-P_{\alpha_{0},1}|>\varepsilon\mid\mathcal{X}_{n})=o_{\mathbb{P}}(1)$ for all $\varepsilon>0$ and all $f\in\mathcal{F}(\alpha_{-},\alpha_{+})$) and that
$d_{K}\big{(}\mathcal{L}(\hat{\mathbb{G}}_{n}^{\ast}(f_{1},f_{2},f_{3})^{\top}\mid\mathcal{X}_{n}),\mathcal{N}_{3}(\bm{0},\Sigma_{G})\big{)}=o_{\mathbb{P}}(1).$ (11.6)
Then, on the complement of the event $\{X_{n,1}^{*}=\dots=X_{n,\omega_{n}}^{*}\}$, the independence Fréchet-log-likelihood ${\bm{\theta}}\mapsto\sum_{i=1}^{\omega_{n}}\ell_{\bm{\theta}}(X_{n,i}^{*})$ has a unit maximizer ${\bm{\theta}}_{n}^{*}=(\hat{\alpha}_{n}^{*},\hat{\sigma}_{n}^{*})$, and that maximizer satisfies
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\hat{\alpha}_{n}\\ \hat{\sigma}_{n}^{*}/\hat{\sigma}_{n}-1\end{pmatrix}=M(\alpha_{0})\hat{\mathbb{G}}_{n}^{*}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1).$ (11.7)
As a consequence, if additionally $\bm{B}_{G}=0$, we have bootstrap consistency in the following sense
$d_{K}\left[\mathcal{L}\left(v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\hat{\alpha}_{n}\\ \hat{\sigma}_{n}^{*}/\hat{\sigma}_{n}-1\end{pmatrix}\,\Big{|}\,\mathcal{X}_{n}\right),\mathcal{L}\left(v_{n}\begin{pmatrix}\hat{\alpha}_{n}-\alpha_{0}\\ \hat{\sigma}_{n}/\sigma_{n}-1\end{pmatrix}\right)\right]=o_{\mathbb{P}}(1).$ (11.8)
Since it concerns empirical means only, the condition in (11.6) is typically a standard result for the bootstrap; see for instance Section 10 and 11 in Kosorok, (2008).
Proof of Theorem 11.7.
First, by Theorem 11.6 (which is Theorem 2.3, Theorem 2.5 and Addendum 2.6 in Bücher and Segers, 2018b, applied to the sample $\mathcal{X}_{n}$), we have
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}-\alpha_{0}\\ \hat{\sigma}_{n}/\sigma_{n}-1\end{pmatrix}=M(\alpha_{0})(\bar{\mathbb{G}}_{n}+B_{n})(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)\rightsquigarrow M(\alpha_{0})\mathcal{N}_{3}(\bm{B}_{G},\Sigma_{G}).$ (11.9)
Next, in view of the fact that (11.6) implies $\hat{\mathbb{G}}_{n}^{*}(f_{1},f_{2},f_{3})^{\top}\rightsquigarrow\mathcal{N}_{3}(\bm{0},\Sigma_{G})$ (unconditionally) by Lemma 2.3 in Bücher and Kojadinovic, (2019), we have
$\mathbb{G}_{n}^{\circ}(f_{1},f_{2},f_{3})^{\top}:=v_{n}\big{(}\hat{\mathbb{P}}_{n}^{\ast}(f_{1},f_{2},f_{3})^{\top}-P_{\alpha_{0},1}(f_{1},f_{2},f_{3})^{\top}\big{)}\rightsquigarrow\mathcal{N}_{3}(\bm{B}_{G},\Sigma_{G}).$
Hence, we may apply Theorem 2.3, Theorem 2.5 and Addendum 2.6 in Bücher and Segers, 2018b to the sample $\mathcal{X}_{n}^{*}$ (unconditionally) to obtain that
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\alpha_{0}\\ \hat{\sigma}_{n}^{\ast}/\sigma_{n}-1\end{pmatrix}=M(\alpha_{0}){\mathbb{G}}_{n}^{\circ}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)$ (11.10)
As a consequence, by (11.9), (11.10) and $\hat{\sigma}_{n}/\sigma_{n}=1+o_{\mathbb{P}}(1)$, we have
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\hat{\alpha}_{n}\\ \hat{\sigma}_{n}^{\ast}/\hat{\sigma}_{n}-1\end{pmatrix}$ $=v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\alpha_{0}-(\hat{\alpha}_{n}-\alpha_{0})\\ \sigma_{n}/\hat{\sigma}_{n}\big{(}\sigma_{n}^{\ast}/\sigma_{n}-1-(\hat{\sigma}_{n}/\sigma_{n}-1)\big{)}\end{pmatrix}$
=$=\{1+o_{\mathbb{P}}(1)\}\,\big{(}M(\alpha_{0})(\mathbb{G}_{n}^{\circ}-\mathbb{G}_{n}-B_{n})(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)\big{)}$
$=M(\alpha_{0})\hat{\mathbb{G}}_{n}^{\ast}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1),$ (11.11)
which is (11.7).
Finally, for the proof of (11.8), let $\mathcal{X}_{n}^{\#}$ denote a second bootstrap sample, generated in the same way as $\mathcal{X}_{n}^{*}$ and independent of $\mathcal{X}_{n}^{*}$, conditionally on $\mathcal{X}_{n}$. Denote the respective estimators and empirical measures/processes by $\hat{\alpha}_{n}^{\#},\hat{\mathbb{P}}_{n}^{\#}$ etc. Then, by the expansions in (11.1),
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\hat{\alpha}_{n}\\ \hat{\sigma}_{n}^{\ast}/\hat{\sigma}_{n}-1\\ \hat{\alpha}_{n}^{\#}-\hat{\alpha}_{n}\\ \hat{\sigma}_{n}^{\#}/\hat{\sigma}_{n}-1\\ \end{pmatrix}=\begin{pmatrix}M(\alpha_{0})(\hat{\mathbb{G}}_{n}^{*})(f_{1},f_{2},f_{3})^{\top}\\ M(\alpha_{0})(\hat{\mathbb{G}}_{n}^{\#})(f_{1},f_{2},f_{3})^{\top}\end{pmatrix}+o_{\mathbb{P}}(1).$
Equation (11.6) and Lemma 2.3 in Bücher and Kojadinovic, (2019) implies that the dominating term on the right-hand side converges weakly (unconditionally) to $\mathcal{N}_{2}(\bm{0},M(\alpha_{0})\Sigma_{G}M(\alpha_{0})^{\top})^{\otimes 2}$, which by another reverse application of that lemma implies
$d_{K}\left[\mathcal{L}\left(v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\hat{\alpha}_{n}\\ \hat{\sigma}_{n}^{*}/\hat{\sigma}_{n}-1\end{pmatrix}\,\Big{|}\,\mathcal{X}_{n}\right),\mathcal{N}_{2}(\bm{0},M(\alpha_{0})\Sigma_{G}M(\alpha_{0})^{\top})\right]=o_{\mathbb{P}}(1).$
The assertion then follows from the triangular inequality and (11.9), noting that $\bm{B}_{G}=\bm{0}$ by assumption. ∎
11.2 Auxiliary results on bootstrap consistency
Lemma 11.8.
Let $\bm{S}_{n},\bm{T}_{n}\in\mathbb{R}^{q}$ with $\bm{S}_{n}\rightsquigarrow\bm{S},\bm{T}_{n}\rightsquigarrow\bm{S}\in\mathbb{R}^{q}.$ Furthermore, let $\mathcal{L}(\bm{S})$ be absolutely continuous with respect to the Lebesgue measure on $\mathbb{R}^{q}.$ We then have
$\lim_{n\to\infty}d_{K}\big{[}\mathcal{L}(\bm{A}_{n}\bm{S}_{n}),\mathcal{L}(\bm{A}_{n}\bm{T}_{n})\big{]}=0,$
for any sequence $(\bm{A}_{n})_{n}\subset\mathbb{R}^{p\times q}$ of matrices.
Proof.
Write $\bm{A}_{n}=(\bm{a}_{n,j}^{\top})_{j=1,\ldots,p}$ with $\bm{a}_{n,j}\in\mathbb{R}^{q}.$ A straightforward argument shows that we may assume $\bm{a}_{n,j}\neq\bm{0}$ for all $j=1,\ldots p$. Next, note that, for all diagonal matrices $\bm{D}=\operatorname{diag}(d_{j})_{j=1,\ldots p}$ with $d_{j}\neq 0$ and all $\mathbb{R}^{p}$-valued random variables $\bm{S},\bm{T}$, we have $d_{K}[\mathcal{L}(\bm{S}),\mathcal{L}(\bm{T})]=d_{K}[\mathcal{L}(\bm{D}\bm{S}),\mathcal{L}(\bm{D}\bm{T})]$. Hence, letting $d_{j}:=\|\bm{a}_{n,j}\|_{2}^{-1}$ and $\tilde{\bm{A}}_{n}=(\tilde{\bm{a}}_{n,j}^{\top})_{j=1,\ldots,p}$ with normed $\tilde{a}_{n,j}=\bm{a}_{n,j}/\|\bm{a}_{n,j}\|_{2}\in\mathbb{R}^{p}$, we have $\tilde{\bm{A}}_{n}=\bm{D}\bm{A}_{n}$ and therefore
$d_{K}\big{[}\mathcal{L}(\bm{A}_{n}\bm{S}_{n}),\mathcal{L}(\bm{A}_{n}\bm{T}_{n})\big{]}=d_{K}\big{[}\mathcal{L}(\tilde{\bm{A}}_{n}\bm{S}_{n}),\mathcal{L}(\tilde{\bm{A}}_{n}\bm{T}_{n})\big{]}.$ (11.12)
Since $\tilde{\bm{A}}_{n}\in[-1,1]^{p\times q}$, the Bolzano-Weierstraß Theorem allows to find a subsequence $n^{\prime}:=n^{\prime}(n)$ such that $\bm{E}=\lim_{n\to\infty}\tilde{\bm{A}}_{n^{\prime}}$ exists. Slutsky’s lemma then yields the weak convergences $\tilde{\bm{A}}_{n^{\prime}}\bm{S}_{n^{\prime}}\rightsquigarrow\bm{E}\bm{S}$ and $\tilde{\bm{A}}_{n^{\prime}}\bm{T}_{n^{\prime}}\rightsquigarrow\bm{E}\bm{S}$. We will show below that the cdf of $\bm{E}\bm{S}$ is continuous. Since the Kolmogorov-metric $d_{K}$ metrizes weak convergence to limits with continuous cdfs (van der Vaart, , 1998, Lemma 2.11), an application of the triangular inequality implies
$d_{K}\big{[}\mathcal{L}(\tilde{\bm{A}}_{n^{\prime}}\bm{S}_{n^{\prime}}),\mathcal{L}(\tilde{\bm{A}}_{n^{\prime}}\bm{T}_{n^{\prime}})\big{]}\leq d_{K}\big{[}\mathcal{L}(\tilde{\bm{A}}_{n^{\prime}}\bm{S}_{n^{\prime}}),\mathcal{L}(\bm{E}\bm{S})\big{]}+d_{K}\big{[}\mathcal{L}(\bm{E}\bm{S}),\mathcal{L}(\tilde{\bm{A}}_{n^{\prime}}\bm{T}_{n^{\prime}})\big{]}=o(1)$
for $n^{\prime}\to\infty$. As the previous argumentation can be repeated for subsequences $n^{\prime}$ of arbitrary subsequences $n^{\prime\prime},$ we then have $d_{K}\big{[}\mathcal{L}(\tilde{\bm{A}}_{n}\bm{S}_{n}),\mathcal{L}(\tilde{\bm{A}}_{n}\bm{T}_{n})\big{]}=o(1)$, which implies the assertion by (11.12).
It remains to show that the cdf of $\bm{E}\bm{S}$ is continuous. For that purpose, by Sklar’s theorem and the fact that copulas are continuous, it is sufficient to show that each marginal cdf of $\bm{E}\bm{S}$ is continuous. Hence, fix $j\in\{1,\dots,p\}$, and note that the $j$th row of $\bm{E}$, say $\bm{e}_{j}^{\top}$, has Euclidean norm 1. We may therefore construct an invertible matrix $\tilde{\bm{E}}=\tilde{\bm{E}}(\bm{e}_{j})\in\mathbb{R}^{q\times q}$ with the first row of $\tilde{\bm{E}}$ being equal to $\bm{e}_{j}^{\top}.$ Then $\tilde{\bm{E}}$ being a diffeomorphism and $\bm{S}$ being absolutely continuous with respect to the Lebesgue-measure on $\mathbb{R}^{q}$ implies the latter for $\tilde{\bm{E}}\bm{S}.$ Hence, for any $y\in\mathbb{R}$, $\mathbb{P}((\bm{E}\bm{S})_{j}=y)=\mathbb{P}\big{(}\tilde{\bm{E}}\bm{S}\in\{y\}\times\mathbb{R}^{q-1}\big{)}=0.$ Since $y\in\mathbb{R}$ was arbitrary, this proves the continuity of $y\mapsto\mathbb{P}((\bm{E}\bm{S})_{j}\leq y).$
Lemma 11.9.
Let $(\Omega,\mathcal{A},\mathbb{P})$ denote a probability space and $p,q\in\mathbb{N}$. For $n\in\mathbb{N},$ let $\bm{X}_{n}\colon\Omega\to\mathcal{X}_{n},\,\bm{W}_{n}\colon\Omega\to\mathcal{W}_{n}$ denote random variables in some measurable space $\mathcal{X}_{n},\mathcal{W}_{n},$ respectively. Let $\bm{S}_{n}=\bm{S}_{n}(\bm{X}_{n})$ and $\bm{S}_{n}^{\ast}=\bm{S}_{n}(\bm{X}_{n},\bm{W}_{n})$ be $\mathbb{R}^{q}$-valued statistics. If
[…]
where $Q$ is absolutely continuous with respect to the Lebesgue-measure on $\mathbb{R}^{q}$ and $d_{w}$ denotes any metric characterizing weak convergence on $\mathbb{R}^{q}$; then
$d_{K}\big{(}\mathcal{L}(\bm{A}_{n}\bm{S}_{n}^{\ast}\mid\bm{X}_{n}),\mathcal{L}(\bm{A}_{n}\bm{S}_{n})\big{)}\xrightarrow{\mathbb{P}}0,\qquad\text{ as }n\to\infty,$
for any sequence $(\bm{A}_{n})_{n}\subset\mathbb{R}^{p\times q}$ of matrices.
Proof.
Let $n^{\prime}\subset\mathbb{N}$ denote a subsequence of $\mathbb{N}.$ By assumption (a), we may choose a further subsequence $n^{\prime\prime}$ of $n^{\prime}$ and $\Omega_{0}\in\mathcal{A}$ with $\mathbb{P}(\Omega_{0})=1$ such that
$\lim_{n\to\infty}d_{w}\big{(}\mathcal{L}(\bm{S}_{n^{\prime\prime}}^{\ast}\mid\bm{X}_{n^{\prime\prime}}),\mathcal{L}(\bm{S}_{n^{\prime\prime}})\big{)}=0\quad\text{ on }\Omega_{0}.$
Hence, since $\bm{S}_{n}\rightsquigarrow Q$ by assumption, we obtain $\mathcal{L}(\bm{S}_{n^{\prime\prime}}^{\ast}\mid\bm{X}_{n^{\prime\prime}})\rightsquigarrow Q$ on $\Omega_{0}.$ Lemma 11.8 then implies
$\lim_{n\to\infty}d_{K}\big{(}\mathcal{L}(\bm{A}_{n^{\prime\prime}}\bm{S}_{n^{\prime\prime}}^{\ast}\mid\bm{X}_{n^{\prime\prime}}),\mathcal{L}(\bm{A}_{n^{\prime\prime}}\bm{S}_{n^{\prime\prime}})\big{)}=0\quad\text{ on }\Omega_{0},$
which lets us conclude. ∎
Appendix A Details on mean estimators in the ARMAX-GPD model
Corollary A.1.
Suppose $(X_{t})_{t\in\mathbb{Z}}$ is an ARMAX-GPD time series as in Model 6.1, for some $\beta\in[0,1)$ and some $\gamma<1/2$. Then, if the block size parameter satisfies $r=o(n)$ and $\log n=o(r^{1/2})$,
$\frac{\sqrt{n/r}}{(r(1-\beta))^{\gamma}}\big{(}\hat{\mu}_{n}^{({\operatorname{mb}})}-\mu_{r}\big{)}\rightsquigarrow\mathcal{N}(0,\sigma^{2}_{{\operatorname{mb}}}),\quad{\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}\},$
with $\sigma^{2}_{{\operatorname{cb}}}=\sigma^{2}_{{\operatorname{sb}}}<\sigma^{2}_{{\operatorname{db}}}$ as provided in (6.2). The ratio $\gamma\mapsto\sigma^{2}_{{\operatorname{db}}}/\sigma^{2}_{{\operatorname{sb}}}(\gamma)$ is presented in Figure 10.
Figure 10: Ratio of the asymptotic variances $\sigma^{2}_{{\operatorname{db}}}/\sigma^{2}_{{\operatorname{sb}}}$ from (6.2).
Proof.
The proof is akin to the one of Equation (4.3) in Bücher and Staud, (2024); most details are omitted for the sake of brevity and we only discuss the calculation of the asymptotic variance parameter $\sigma^{2}_{\operatorname{sb}}$ (note that $\sigma^{2}_{\operatorname{db}}$ simply corresponds to the variance of the $\mathrm{GEV}(\gamma)$-distribution by definition).
We start by considering the case $\gamma\neq 0.$ For $i\in\{1,2\}$, the random variables $S_{i,\xi}:=(1+\gamma Z_{i,\xi})^{1/\gamma}$ are standard exponentially distributed, and we have $\operatorname{Cov}(Z_{1,\xi},Z_{2,\xi})=\gamma^{-2}\operatorname{Cov}(S_{1,\xi}^{-\gamma},S_{2,\xi}^{-\gamma})=:\gamma^{-2}C_{\xi}.$ For $\gamma<0,$ Equation (C.8) in the supplement of Bücher and Staud, (2024) implies that
$\int_{0}^{1}C_{\xi}\,\mathrm{d}\xi=2\gamma^{2}\Gamma(-2\gamma)\int_{0}^{1/2}(\alpha_{2\gamma}(w)-1)\big{(}w^{-\gamma-1}(1-w)^{-\gamma-1}\,\mathrm{d}w)=2\gamma^{2}\Gamma(-2\gamma)I(\gamma),$
hence $\sigma^{2}_{{\operatorname{sb}}}=4\Gamma(-2\gamma)I(\gamma).$ For $\gamma>0,$ the display above (C.9) in the same reference implies
$\int_{0}^{1}C_{\xi}\,\mathrm{d}\xi=-\gamma\frac{\Gamma(1-2\gamma}{2}I(\gamma).$
Hence, $\sigma^{2}_{{\operatorname{sb}}}=-\Gamma(1-2\gamma)/\gamma I(\gamma).$
Finally, for $\gamma=0$ consider the transformed random variables $S_{i,\xi}=\exp(-Z_{i,\xi})$ for $i\in\{1,2\}$, which gives $\operatorname{Cov}(Z_{1,\xi},Z_{2,\xi})=\operatorname{Cov}(\log S_{1,\xi},\log S_{2,\xi})=C_{\xi}.$ By formula (C.11) from the same reference we have
$\int_{0}^{1}C_{\xi}\,\mathrm{d}\xi=\int_{0}^{1}\frac{1}{w(1-w)}\int_{0}^{1}-\log(A_{\xi}(w))\,\mathrm{d}\xi\,\mathrm{d}w,$
where $A_{\xi}(w)=\xi+(1-\xi)\big{(}w\wedge(1-w)\big{)}$ is the Pickands-dependence function of the associated Copula $C_{\xi}$ of the Marshall-Olkin distribution with dependence parameter $\xi.$ It follows that
$\int_{0}^{1}C_{\xi}\,\mathrm{d}\xi=2\int_{1/2}^{1}\frac{1}{w(1-w)}\int_{0}^{1}-\log(\xi+(1-\xi)w)\,\mathrm{d}\xi\,\mathrm{d}w=\log 4-1,$
which implies the asserted formula. ∎
Appendix B Runtime comparison
In classical situations, the computational cost of the bootstrap depends linearly on the number of bootstrap replications and is therefore high if a single evaluation of a statistic of interest is computationally intensive. Since both the sliding and the circular block maxima samples are much larger than the plain disjoint block maxima sample (sample sizes $n$ vs. $n/r$, respectively), one may naively think that the former methods require substantially more computational resources. As we will argue below and prove with simulations, this naive heuristic is not correct.
For any block maxima method, the starting point for a single evaluation of a statistic of interest is the calculation of the respective block maxima samples, which requires evaluating $O(n/r)$ maxima for the disjoint block maxima method, or $O(n)$ maxima for the sliding and circular block maxima method. Subsequently, when the statistic of interest is applied to one of the samples, the fact that both the sliding and the circmax samples can be efficiently stored as a weighted sample of size $O(n/r)$ implies that the additional computational cost of the latter two methods is, approximately, only a constant multiple of the additional cost for the plain disjoint block maxima method. Next, for the bootstrap approaches proposed within this paper, no additional evaluation of maxima is ever required, whence, overall, the only major difference between the three approaches is the initial calculation of the block maxima samples. Therefore, the relative computational effort should only depend on $n$, and only moderately.
The above heuristic has been confirmed by Monte Carlo simulations. Exemplary results are presented in Figure 11, which rely on simulated data from Model 6.2 (parameters: $\beta=0.5$ and $\alpha=1.5$) with fixed block size $r=90$ and total sample size ranging form $40\cdot 90=3,600$ up to $100\cdot 90=9,000$. The target parameter is the runtime for calculating $B\in\{250,500,750,1,000\}$ bootstrap replicates of $\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}$ from (5.1) for ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}(2),{\operatorname{cb}}(3)\}$, assessed by taking the median over $N=500$ repetitions each. The disjoint blocks method has been considered as a benchmark, whence we depict relative runtimes with respect to that method. We find that, as expected, the relative runtime is mostly depending on $n$, with only a moderate loss in performance for the circmax-method. Similar results were obtained for other estimators and models.
Figure 11: Relative median runtimes of different bootstrap algorithms for bootstrapping $\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}$ (relative to the runtime of the disjoint blocks bootstrap) for fixed blocksize $r=90$ as a function of the effective sample size and for different numbers of bootstrap replicates $B.$
Acknowledgements
Financial support by the German Research Foundation (DFG grant number 465665892) and by Ruhr University Research School (funded by Germany’s Excellence Initiative - DFG GSC 98/3) is gratefully acknowledged. Computational infrastructure and support were provided by the Centre for Information and Media Technology at Heinrich Heine University Düsseldorf. The authors are grateful to Johan Segers for fruitful discussions on the circular block maxima method, and to the participants of the Oberwolfach Workshop on “Mathematics, Statistics, and Geometry of Extreme Events in High Dimensions” for their valuable comments.